The ACPC Network Guide Version 1.1

The ACPC Network Group

consisting of

H. W. Loidl, RISC-Linz (Coordinator)

H. Wabnig, University Vienna J. Fritscher, TU-Vienna S. Grabner, GUP-Linz

E. Cikan & P. Kutschera, University Vienna N. Ulamec & M. Salchegger, RIST++ Salzburg

July 6, 1993

Contents

1 Introduction 2 1.1 What is the ACPC? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Aims of the ACPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 The ACPC Network . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Guest Accounts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 The Account Application Procedure . . . . . . . . . . . . . . . . . . . . 5 1.6 The Official ACPC Anonymous FTP Server . . . . . . . . . . . . . . . . 8 1.7 The ACPC Mailing List . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2 Network Connections 10

3 The Groups in the ACPC 16

3.1 RISC-Linz — Research Institute for Symbolic Computation, Johannes

Kepler University . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.2 Institute for Applied Computer Science and Information Systems, Uni-

versity of Vienna . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.3 GUP-Linz — Department for Computer Graphics and Parallel Proces-

sing, Institute of Computer Science, Johannes Kepler University Linz . . 18 3.4 Institute for Software Technology and Parallel Systems, University of

Vienna (PPS — Programming Languages, Parallel Systems) . . . . . . . 19 3.5 RIST++ — Research Institute for Software Technology + Paris Lodron

University Salzburg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.6 List of Documentation that is available at each ACPC Group . . . . . . 20

3.6.1 RISC-Linz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.6.2 Institute for Applied Computer Science and Information Sy-

stems, University of Vienna . . . . . . . . . . . . . . . . . . . . . 21 3.6.3 TU-Vienna . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.6.4 GUP-Linz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4 Facilities in the Network 23

4.1 Group Description of RISC-Linz . . . . . . . . . . . . . . . . . . . . . . 23 4.1.1 Sequent Symmetry S81 . . . . . . . . . . . . . . . . . . . . . . . 24 4.1.2 Transputer Network . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.1.3 Silicon Graphics Iris VGX Graphics Supercomputer . . . . . . . 53

4.2 Group Description of the Institute for Applied Computer Science and Information Systems, University of Vienna . . . . . . . . . . . . . . . . . 54 4.2.1 MultiCluster-3DE . . . . . . . . . . . . . . . . . . . . . . . . . . 55

4.3 Group Description of TU-Vienna . . . . . . . . . . . . . . . . . . . . . . 58 4.3.1 Sequent Balance B21 . . . . . . . . . . . . . . . . . . . . . . . . . 58 4.3.2 Intel Hypercube iPSC/860 . . . . . . . . . . . . . . . . . . . . . . 61

4.4 Group Description of the Theoretical Chemistry Group . . . . . . . . . 64 4.5 Group Description of GUP-Linz . . . . . . . . . . . . . . . . . . . . . . . 65

4.5.1 nCUBE 2 Model 1 . . . . . . . . . . . . . . . . . . . . . . . . . . 65 4.6 Group Description of the Institute for Software Technology and Parallel

Systems, University of Vienna (PPS) . . . . . . . . . . . . . . . . . . . . 69 4.6.1 Intel iPSC/860 Hypercube . . . . . . . . . . . . . . . . . . . . . . 70 4.6.2 Transputersystem . . . . . . . . . . . . . . . . . . . . . . . . . . 72 4.6.3 MasPar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

4.7 Group Description of RIST++ Salzburg . . . . . . . . . . . . . . . . . . 78 4.7.1 SUN 3/60 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 4.7.2 SPARC 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 4.7.3 SPARC ELC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 4.7.4 CONVEX C220 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 4.7.5 ALPHA Cluster (6+1 DEC 3000 Alpha AXP 400) . . . . . . . . 83

A Table of All Available Parallel Machines in the ACPC Network 85

B Alphabetical List of Persons in the ACPC 86

Abstract

The ACPC Network Guide is a manual that gives the reader an overview of the machines in the ACPC Network and the software that is installed on these machines. Therefore, it gives the user of one of these machines the most important information that is needed for the work on this machine and it refers to additional manuals for more complex software packages where this is necessary.

However, before this overview is given, in the three introductory chapters the structure and the aims of the ACPC are discussed. The first chapters also contain some general information about the ACPC Network that might also be of interest for people who do not intend to work on one of the machines in this network. A reference to an FTP server, containing information about the ACPC Network and the ACPC in general is given. The ACPC Mailing List, which is used for distributing information among the groups in the ACPC, is also mentioned in this general part of the manual. Furthermore, this part also contains a short description of the groups in the ACPC and of their research interests in the area of parallel computation.

The main part of the manual is then formed by one section for each ACPC working group that contains some information about the group and its members, a description of the hardware at this group and a description of the software that is installed on these machines.

The appendices summarize information that is already present but spread around in the other chapters. As we think that the information presented there will be needed quite frequently, the appendices can be seen as a kind of quick reference.

About the Authors

The authors of this manual are the members of the ACPC Network Group. Each working group in the ACPC has one representative, the Network Coordinator of that group, who is responsible for keeping the information about his working group in this manual up to date. The Network Coordinator is also the contact person for people who want to work on a machine of the group he is representing or just for getting more information about the machines or the software on the machines of that group.

As one important task of the ACPC Network Group is to keep this manual up to date, one member of this group has to coordinate the changes that have to be made. This coordinator of ‘The ACPC Network Guide’ also takes care of the general parts and of the appendices in this manual. This coordinator is currently H. W. Loidl from RISC-Linz.

Currently, the ACPC Network Group consists of the following people:

Network Coordinator  ACPC Group  Email / Telephone
H. W. Loidl  RISC-Linz  hwloidl@risc.uni-linz.ac.at

07236 3231 66

J. Fritscher TU-Vienna, fritscher@edvz.tuwien.ac.at
Theoretical Chemistry  0222 58801 5505

Group (Vienna)

S. Grabner GUP-Linz grabner@gup.uni-linz.ac.at

0732 2468 884

E. Cikan University Vienna cikan@par.univie.ac.at
(PPS)  0222 392647 201
P. Kutschera University Vienna peter@par.univie.ac.at
(PPS)  0222 40103 2789
H. Wabnig University Vienna wabnig@ani.univie.ac.at

0222 4086366 17

N. Ulamec RIST++ Salzburg ulamec@kirk.coma.sbg.ac.at

0662 8044 6770

M. Salchegger  RIST++ Salzburg salchegg@spock.coma.sbg.ac.at

0662 8044 6771

Chapter 1

Introduction

1.1 What is the ACPC?

The Austrian Center for Parallel Computation (ACPC) is a joint organization formed by seven institutes at the universities of Linz, Salzburg and Vienna. The ACPC was founded in 1989 to promote research, education and industrial cooperation in the area of parallel computation in Austria and is supported by the Austrian Research Foundation and the Austrian Ministry of Science and Research. At present, approximately 50 researchers at the levels of professor, senior and junior researcher, and doctoral student from these seven groups are united in the ACPC effort.

Each group of the ACPC is specialized in a different subarea of parallel computation (parallel symbolic computation, parallel numerical applications, parallelization of programming languages, parallel computer graphics and visualization, theory of parallel computation, parallel hardware and software systems, …). Hence, the ACPC groups profit from a lot of synergy effects by several joint research projects.

Participants in the ACPC are at present seven academic groups headed by (in alphabetical order):

Prof. Bruno Buchberger, RISC-Linz

Prof. Günter Haring, Institute for Applied Computer Science and Information Systems, University of Vienna

Dr. Wolfgang Kleinert/Doz. Christoph Überhuber, Technical University of Vienna

Prof. Peter Schuster/Doz. Othmar Steinhauser, Theoretical Chemistry Group, University of Vienna

Prof. Jens Volkert (Spokesman), Department for Computer Graphics and Parallel Processing, Institute for Computer Science, University of Linz Prof. Hans P. Zima, Institute for Software Technology and Parallel Systems, University of Vienna

Prof. Peter Zinterhof, RIST++ Salzburg

1.2 Aims of the ACPC

The ACPC aims at focusing and coordinating the Austrian academic research and graduate education in the area of parallel computation. It starts from the past and present academic research activities at the major Austrian universities and organizes a series of interrelated new research projects in the area of software for parallel architectures and applications of parallel computation. The ACPC started a graduate special curriculum in parallel computation as a part of a computer science or mathematics master’s or Ph.D. program. The ACPC puts a particular emphasis on motivating Austrian software and hardware companies to cooperate in pursuing the goals of the ACPC and in initiating projects that transfer foundational research in parallel computation into industrial products (software, systems).

Summarizing, the main activities for achieving these goals of the ACPC are:

• The ACPC organizes joint research projects between the various members of the ACPC with a special view to achieving synergy. Many of these projects also include partners from outside the ACPC, thereby embedding the ACPC in the international activities in the research in parallel computation.

• The ACPC has already designed a joint curriculum for diploma and doctoral students who want to specialize in the area of parallel computation so that students can easily get credit for courses taken at a different site of the ACPC.

• The ACPC cooperates with several industrial partners in joint work leading to a technological transfer and marketable products. Furthermore, the new results of ongoing projects are offered and presented to the industry such that immediate profit can be taken out of these scientific developments.

• The ACPC encourages researchers and students to visit and to study at different sites within the ACPC. By this, the researchers and the students can get profit out of the research in different subareas of parallel computation that is done at the different sites.

• To improve the knowledge transfer in the area of parallel computation not only inside the ACPC, which is done by regular internal ACPC workshops, but also with other institutes working in this area, the ACPC has organized the ‘1^{st} International Conference of the Austrian Center for Parallel Computation’ in Salzburg, in October 1991. This was the first of a series of conferences that will be held every two years. Furthermore, the ACPC regularly organizes workshops that deal with special aspects of parallel computation. Another important means for the international knowledge transfer is the ‘ACPC Technical Report Series’ in which all the groups of the ACPC publish their new research results.

• The ACPC has set up a powerful network over which all partners of the ACPC can access all parallel machines available at the various sites. Therefore, all researchers in the ACPC have access to almost all modern parallel architectures. This manual aims at giving the reader an overview of these parallel architectures their characteristics, location in the network and the software that is installed on these machines.

1.3 The ACPC Network

The ACPC Network is a network of parallel computers located at the individual institutions that form the ACPC. In order to reach each of these computers they have been combined to the ACPC Network. Therefore, it should be possible for every person from one of the ACPC groups (but also from outside) to reach any computer in the network. Thereby, the ACPC Network grants access to essentially every interesting parallel architecture.

From the hardware side the network is based on the already existing connections that combine the universities in Austria. As the whole network is itself a part of the worldwide Internet, the machines can be reached from any part of the world that has access to the Internet. For a more detailed description of the hardware upon which the network is based see Chapter 2.

However, just building up the hardware and operating system layer for combining various machines in a network is of course not enough to give other users the opportunity of working on remote machines. For this purpose, also a description of the software and of the working environment on the machines in the network is necessary. Therefore, the main part of this manual (Chapter 4) deals with the description of the hardware and of the software that is available at each ACPC group. This should be used as a first information to get a quick overview of the available facilities in the network. However, if somebody is interested in one special software product that is installed on one of the machines he/she may refer to the documentation of the software products that is available at each of the ACPC groups^{1}. However, that part of the documentation that has been distributed over all groups has been kept as small as possible. Hence, one should also contact the network coordinator of the group where the software product has been installed to get more information.

1.4 Guest Accounts

In the ACPC Network, on each machine some guest accounts have been installed. Hence, each ACPC group has one such guest account on each machine in the network^{2}.

1For a list of the documentation that is available at each group see Section 3.6.

2See Section 4 for a description of each of the machines in the ACPC Network.

The purpose of this account is to offer all members of the group the possibility to have a look on each available machine, in order to decide whether it is interesting to apply this particular machine. This opportunity is especially intended for students who want to get an overview of the available hardware without themselves having to visit all the institutes of the ACPC. As there are now many different architectures available in the ACPC Network, every person working in the ACPC can choose the most appropriate architecture for his application.

However, these guest accounts are only intended to get a very first experience. Typical actions performed under this account will be for example the reading of on-line manuals to study some details or just having a look what is new on the machine. As soon as somebody wants to do some real work this person should apply for a personal account and then work under this account. For an explanation of this application procedure see Section 1.5.

As such guest accounts are always a potential security hole for each machine, there may be some security measures taken to avoid an unauthorized usage. However, there is no global security strategy but each system administrator may decide about this independently. The user of a guest account must therefore be prepared to be confronted with one or several of the following security measures:

• The user might be asked for his email address such that the system gets information who is actually logged in.

• The permissions of some files like .cshrc might be set such that the user can not modify them.

• The group of the user may be set in such a way that it will not allow him any additional actions.

• The password of the guest account might have to be changed regularly.

These security measures may lead to a rather inconvenient guest account. But as this account is only intended for some simple actions like listed above this should be no problem. Hence, the general rule is that whenever some serious work is done on a machine, a personal account shall be installed.

1.5 The Account Application Procedure

This section will describe how to get an account on any machine in the ACPC Network. The main motivation for introducing this standard application procedure was to get the information about who is actually working on a certain machine. We have tried to make the whole procedure as convenient for the user as possible. However, some minimal effort in applying for an account is inevitable.

All you have to do for getting an account on a certain machine is to fill in the application form that is shown in Figures 1.1 and 1.2 and to send the form to the system

Autrian Center for Parallel Computation

Account Application Form

This application is to be filled out for an account on a computer system managed by the ACPC. System Name (as described in the ACPC Manual): Name: Organization: Telephone: E-mail address: Mail address: My intended research on this system is:

Project Title:

Are you a member of the ACPC? Do you have any special requirements (hardware, software or other resources)?

Figure 1.1: Application Form (Part 1)

We must request you to comply with the following guidelines when using your account on the above mentioned computer system. If you do not, we are obliged to safeguard the interests of other users by removing your account. Should your project require substantial machine resources, special approval must be given. Please indicate this in the description of your intended research. You will be notified by e-mail when you account is installed. The machine may go down for experimental work or for systems operations with little notice. The disk is not to be used for backup purposes. Due credit will be given in publications reporting on work done on this machine. A copy of each such publication will be sent to the system managing site (according to the ACPC Manual) unless it is published in the ACPC Technical Report series. I agree to abide by the guidelines for use of the system as set forth above and any further rules on machine usage should they become necessary. I understand that failure to comply with any regulations will result in the removal of my account. Name: Signature: Date: Group leader: Signature:

Figure 1.2: Application Form (Part 2) administrator of that machine. As soon as the account has been created, the user will be informed about the login name and the password via email. The application form that is shown in Figures 1.1 and 1.2 is available at all ACPC groups and as a file at the official ACPC anonymous FTP server. The filled in form can be sent via FAX to speed up the whole process of account application.

1.6 The Official ACPC Anonymous FTP Server

In the ACPC the exchange of files is usually handled in such a way that the sender puts files on a special FTP server, the so called official ACPC anonymous FTP server, and announces the availability of this file to the receiver. This procedure is especially recommended if the file has to be sent to several persons or if it might be of interest to several persons or if it should be made publicly available for a longer time.

The address of the official ACPC anonymous FTP server is: ftp.risc.uni-linz.ac.at directory pub/acpc. Physically, the FTP server is located at RISC-Linz in Hagenberg.

The information in the FTP server is structured into several subdirectories:

• The directory Manual contains the sources of the ACPC Manual as well as the complete current version of the manual in the files acpc-manual.dvi.Z and acpc-manual.ps.Z. The latter two files are a compressed dvi and a compressed Postscript file, respectively. This directory is further structured into subdirectories:

– For each group there is one subdirectory containing the information that is specific to this group (this includes the information in Chapter 3 and Chapter 4).

– The directories general and appendices contain general information that appears in the first chapters of the manual and in the appendices of the manual, respectively.

– The directory forms contains the current hardware and software forms that must be used to incorporate new hardware or software into the manual. Furthermore, this directory contains the account application form that is also shown in Figures 1.1 and Figures 1.2. The other files in the forms directory specify the structure of the overall manual and of the group specific information.

– The directory suggest contains suggestions for changes in the ACPC Manual.

However, this information should be only of interest for the ACPC Network Group, who maintains the manual. For others only the above two files, containing the manual itself will be relevant.

• The directory Misc contains miscellaneous information not fitting elsewhere.

• The file News contains important news that might be of interest for all ACPC groups.

• There are subdirectories for several ACPC groups that are maintained by these groups and that contain informations of these groups that has to be published or sent to other people.

As an alternative to using the ACPC anonymous FTP server for sending files normal email should be used if the files should be sent only once to one person. For this purpose we recommend the following procedure:

1. Make a tar file out of all files that have to be sent. This file should have the extension ‘.tar’ (e.g. foo.tar).

2. Compress this file using standard compress. This is not only important to minimize the data that has to be actually sent but it is also important as some mailers might restrict the size of emails they are sending. Using GNU Zip instead of compress will yield even shorter files. However, if you want to use this compression format please make sure that the receiver also has the GNU Zip program available. The resulting file should have extension ‘.Z’ if it has been compressed or ‘.z’ if it has been zipped (e.g. foo.tar.Z).

3. As the whole email system is intended to send ASCII text files only, there will be problems with sending binary files (such as dvi files). Certain bytes in these files will be interpreted (such as CTRL-D as end of file) and therefore the files in general won’t be received in the same form as they have been sent. Therefore, all binary files must be uuencoded before they can be sent (using the command uuencode <file> <name>). The resulting file should have extension ‘.uue’ (e.g. foo.tar.Z.uue). This file now can be sent via email.

1.7 The ACPC Mailing List

In order to facilitate and to speed up the communication among all members of the ACPC the ACPC mailing list has been created. The idea of this mailing list is that all emails being directed to the mailing list are automatically sent to all subscribers of the mailing list. Therefore, any message can easily be sent to all interested persons in the ACPC. However, as every email that is sent to the ACPC mailing list will be received by everybody in the ACPC, this should be used only for topics that really affect all ACPC groups.

The email address of the ACPC mailing list is: ACPC@aearn.bitnet.

It has to be emphasized that sending to the ACPC Mailing List is not a privilege of ACPC members only. Especially, it can be used for the announcement of workshops, talks or other events in the area of parallel computation.

Network Connections

In this chapter the network connections between the ACPC groups will be discussed. Some basic knowledge about the capacity of these hardware connections is very important to understand how the interactive work on a remote machine will look like.

First of all, the capacities of the long range connections are stated:

• Vienna – Linz^{1}: 128 kb/s.

• Linz – Hagenberg: 64 kb/s.

• Linz – Salzburg: 64 kb/s.

• Vienna – Salzburg: 64 kb/s.

Apart from the direct connections between the institutes there is also a connection between Linz and Graz. This connection is important, as the logical connection between Linz and Vienna may be routed over this connection^{2}. However, currently the routing between Linz and Vienna is done dynamically either over the direct connection or over Graz, depending on how much load is on each of these connections. Therefore, it can’t be said beforehand which way a packet from Linz to Vienna will take. On various machines in the ACPC Network a tool called traceroute is installed that gives, when it is invoked with the target machine as its parameter, the way of the connection to the target machine by printing each intermediate machine or router. It is strongly recommended to use this tool (if available) if there are problems with the connection between two machines in the ACPC Network.

The capacities of the connections from and to Graz are:

• Vienna – Graz: 128 kb/s.

• Linz – Graz: 64 kb/s.

CHAPTER 2. NETWORK CONNECTIONS 11 Hagenberg Sequent Symmetry Transputer Silicon Graphics RISC-Linz Linz

nCUBE 2 Salzburg

{Convex \atop C220}

Alpha Cluster Intel Hypercube iPSC/860 Sequent Balance TU-Vienna

Vienna Uni-Vienna

RIST++ GUP-Linz Intel Hypercube iPSC/860 MasPar MP- 1101A

PPS-Vienna

64 kb/s 64 kb/s 128 kb/s Transputer

MC-3DE Transpu- ter

[figure] [figure] [figure] [figure] [figure]

Vienna

Vienna

64 kb/s 64 kb/s

128 kb/s

64 kb/s

Graz

Figure 2.1: ‘Map’ of the ACPC Network

RISC

RISC-Linz Sequent Symmetry 5/6/17 (0 %)
Transputer  7/10/12 (0 %)
Uni-Vienna  MC-3/DE Transputer 125/306/767 (0 %)
TU-Vienna Sequent Balance 113/728/4027 (0 %)
Intel Hypercube  102/482/2864 (0 %)
GUP-Linz  nCube2  64/67/69 (0 %)
Uni-Vienna  MasPar  143/554/3400 (0 %)
(PPS)  Intel Hypercube  150/692/4340 (0 %)
Transputer  99/342/858 (0 %)
RIST++ Salzburg Convex  —
Alpha Cluster  107/143/389 (0 %)

Table 2.1: Ping Times of the ACPC Machines from RISC-Linz

To give the reader a global overview of the whole ACPC Network, Figure 2.1 shows a ‘map’ of this network containing all machines currently available. The quarter arc between the lines to Vienna and to Graz originating in Linz denotes the dynamic routing over the connection with the smallest load. Furthermore, the capacities of the long range connections are also shown in this figure. For a more detailed information on the various computers in the network see Chapter 4.

However, for an exact characterization of the speed of a connection the information about its theoretical capacity is not sufficient. Better information is provided by the ping command that allows the user to check the time required for sending a packet of a certain size to another machine. Some tests with this command have been performed on most of the machines in the ACPC Network. The results of these tests can be seen in Figures 2.1 to 2.5.

However, these times must not be taken as absolute values for the general quality of a certain connection! It must be emphasized that these times may vary very much according to the daytime when it is tested. But these timings should give the reader a rough feeling how fast a certain connection would be.

Experience Reports

The rest of this chapter contains some experience reports about the work on remote machines.

1This connection will be upgraded to 256 kb/s at the end of May 1993.

2Note that this also affects the machines in Hagenberg and probably those in Salzburg when they set up a connection to a machine in Vienna.

CHAPTER 2. NETWORK CONNECTIONS 13

Uni-Vienna

RISC-Linz Sequent Symmetry 124/320/700 (0 %)
Transputer  266/961/1444 (0 %)
Uni-Vienna  MC-3/DE Transputer  local
TU-Vienna  Sequent Balance  37/54/161 (0 %)
Intel Hypercube  34/53/231 (0 %)
GUP-Linz  nCube2  332/588/1210 (0 %)
Uni-Vienna  MasPar  77/77/80 (0 %)
(PPS)  Intel Hypercube  80/81/98 (0 %)
Transputer  30/30/35 (0 %)
RIST++ Salzburg Convex  —
Alpha Cluster  121/791/4979 (43 %)

Table 2.2: Ping Times of the ACPC Machines from the University of Vienna

TU-Vienna

RISC-Linz Sequent Symmetry 100/407/1421
Transputer  —
Uni-Vienna  MC-3/DE Transputer  29/30/42
TU-Vienna  Sequent Balance  7/10/14
Intel Hypercube  3/3/5
GUP-Linz  nCube2  50/1258/5364
Uni-Vienna  MasPar  48/58/139
(PPS)  Intel Hypercube  —
RIST++ Salzburg Convex  —
Alpha Cluster  —

Table 2.3: Ping Times of the ACPC Machines from the TU-Vienna CHAPTER 2. NETWORK CONNECTIONS 14

GUP-Linz

RISC-Linz Sequent Symmetry 62/112/685
Transputer  61/84/206
Uni-Vienna  MC-3/DE Transputer 83/952/6389 (7 % loss)
TU-Vienna Sequent Balance 109/305/1956
Intel Hypercube  60/199/619
GUP-Linz  nCube2  1/1/2
Uni-Vienna  MasPar  121/504/1234
(PPS)  Intel Hypercube  116/600/1843
Transputer  55/125/251
RIST++ Salzburg Convex  —
Alpha Cluster  —

Table 2.4: Ping Times of the ACPC Machines from GUP-Linz

Uni-Salzburg

RISC-Linz Sequent Symmetry 103/158/150 (0 %)
Transputer  —
Uni-Vienna  MC-3/DE Transputer  —
TU-Vienna  Sequent Balance  101/322/691 (0 %)
Intel Hypercube  84/269/609 (0 %)
GUP-Linz  nCube2  43/121/366 (0 %)
Uni-Vienna  MasPar  —
(PPS)  Intel Hypercube  120/462/1019 (0 %)
RIST++ Salzburg Convex  0/0/0 (0 %)
Alpha Cluster  0/0/0 (0 %)

Table 2.5: Ping Times of the ACPC Machines from the University of Salzburg Connection RISC – University Vienna

The interactive work on this connection is quite slow. At least it is significantly slower than on the connection between RISC and the TU-Vienna. Hence, if some serious work should be done over this connection it is recommended to do the editing of files locally and then to transfer the files to the remote computer. This style of work may be a little inconvenient but as soon as one gets accustomed to this way of working there should be no bigger problems with it.

Connection RISC – TU-Vienna

This connection is faster than the connection to the University Vienna. This can also be seen from Table 2.3. With the speed on this connection it is very well possible to do a remote editing of files. Although, there will be a little delay this should not hinder the work.

Connection RISC – University Linz

This connection is the fastest one from RISC within the ACPC Network. If one works from RISC remotely on a machine at the University Linz there is hardly any delay recognizable. Therefore, the work could be done as if one would work locally.

The Groups in the ACPC

This chapter contains a short introduction of each of the ACPC groups. By this the reader should get some information on what parts of parallel computation each of the groups is mainly interested in.

This chapter should only give a first impression of the areas in which the different groups are working. For further information (e.g. about ongoing projects) the interested reader should contact the several groups directly.

3.1 RISC-Linz — Research Institute for Symbolic Computation, Johannes Kepler University

At RISC-Linz, whose chair is Prof. Bruno Buchberger and which is located in the castle of Hagenberg near Linz, Parallel Computation is handled in close connection with Symbolic Computation. Symbolic Computation is that part of computer science respectively mathematics that is dedicated to the solution of problems with non-numerical (“symbolic”) objects. It serves as the basis for many advanced software topics, such as

• computer aided design and geometric modelling,

• robotics and softautomation,

• computer algebra systems,

• expert systems and artificial intelligence,

• logic and functional programming,

• scientific software.

In all these areas, there exist many very interesting and powerful, but also timeconsuming, algorithms. Thus, many problems become intractable very soon when handled on conventional sequential computer architectures. Therefore, Parallel Computation has become an integrated part of Symbolic Computation.

The research in Parallel Computation at RISC-Linz has its roots in 1978, when a parallel architecture and programming language for Symbolic Computation were created. Today, the Parallel Computation Group at RISC-Linz consists of two faculty members, four PhD students and several diploma students. The main research interests of this group are:

• the design and implementation of parallel algorithms for computer algebra, computational logic and computer geometry,

• the design and implementation of parallel software systems for implementing these algorithms, and

• the parallelization of declarative (functional and logic) programming languages.

A special emphasis is put on the synergy between these research areas in Parallel Computation and also between the work in other areas of Symbolic Computation and the work in Parallel Computation.

3.2 Institute for Applied Computer Science and Information Systems, University of Vienna

The main research topics in the parallel processing field are:

• performance evaluation of multiprocessor systems

• process to processor mapping and process scheduling

• development environments and methodology for parallel software

• parallel and distributed simulation

• neural network models and their simulation on parallel hardware

The term performance evaluation covers all aspects of using multiprocessor hardware to gain efficiency. A special focus is on performance modelling based on timed Petri nets (deterministic and stochastic). Such models are used to characterize the factors influencing the performance of working parallel systems (e.g. algorithms for parallel programs, hardware descriptions, process to processor mapping etc.) and are solved either analytically or by simulation. Performance indices derived by the model evaluation are used to analyze the efficiency of parallel programs, the utilization of resources (hardware) and last but not least provide the basis for process to processor mappings. Process to processor mapping is a key problem in parallel processing which cannot be treated analytically in its instance as a graph matching problem because of NPcompleteness. A lot of assumptions in current approaches to solve the problem turn out to be very weak, whereas a lot of characteristics of both the parallel program (like synchronization) and the parallel hardware are not represented in the mapping models. Simulated annealing as a general technique in combinatorial optimization offers a promising facility to the practical treatment of (small) practical problem instances.

Parallel programming techniques (divide-and-conquer, compute-aggregate-broadcast, pipelining, etc.) should be provided to the developer in a rather automated manner. They should be a part of a general parallel program development environment, containing tools for (advanced) editing, compilation, debugging, parallelization and code optimization, performance prediction and monitoring, visualization etc. Usability and efficiency of such tools depend highly on the user interface and the modes of interaction. Both these aspects, tool-support and sophisticated user interfaces, are becoming more and more important with the upcoming multiprocessor workstation technology. The availability of a performance oriented parallel program development methodology is quite attractive from a scientific as well as from an economic point of view.

Today the neural processing approach is successfully operating in constraint satisfaction, pattern association, content-addressable memory, control, motion, vision, coordination applications etc. Special purpose neural networks have shown the power of the non-algorithmic processing methodology encountering qualities like distributed knowledge representation, tendential behaviour, learning, generalization, fault tolerance etc., and beating traditional AI methodologies in several cases. The massive parallelism offered by neural network simulation models can naturally often be exploited, and parallel execution can speed up simulations to bring them into practical use. The usage of parallel processing hardware promises not only the quantitative improvement of reducing simulation times, but gives also a perspective for a qualitative improvement of neural network models.

3.3 GUP-Linz — Department for Computer Graphics and Parallel Processing, Institute of Computer Science, Johannes Kepler University Linz

At GUP-Linz (whose chair is Prof. Jens Volkert), research in parallel computation began in late 1990. This department consists of Prof. Volkert, three Ph.D. students and several diploma students. The main research interests of this group are:

• Monitoring & Debugging of Parallel Programs (especially for distributed memory machines)

• Visualization of Data Partitioning and Graphical Domain Decomposition for Distributed (or Virtual Shared) Memory Machines • Visualization of Multidimensional Data (in parallel)

• Hardware Modelling for Performance Analysis of Parallel Programs

The main basis for this work is a nCUBE 2 distributed memory hypercube with location at the Computing Center of the University Linz.

3.4 Institute for Software Technology and Parallel Systems, University of Vienna (PPS — Programming Languages, Parallel Systems)

One of the most fundamental challenges facing computer science today is the need to develop the algorithms and programming tools required to harness the vast computing power of modern high-performance computers. The commercially available machines that fall into this (roughly defined) class today include vector and array computers, parallel architectures with a large number of processors, and networks of high-performance workstations. These machines provide their users with a peak performance of up to 10 billion floating point operations per second, thus opening the way for the solution of problems which seemed computationally quite out of reach until recently, and spawning interest in a whole range of new problems.

High-performance computers have become a fundamental tool of science. While computers have long been used to organize and control scientific experiments and analyze their results, the new machines can do much more than that: they play an important role in theoretical science too, enabling scientists to construct and test models in their search for theories to describe natural phenomena.

Our interest focusses on programming environments for high-performance computers, in particular on the following areas:

• automatic parallelization

• performance analysis

• advanced compiling techniques

• use of knowledge-based methods for the intelligent support of program transformation strategies

• intelligent man-machine interfaces

We believe that these areas are of vital importance if high-performance computers are to be effectively utilized.

3.5 RIST++ — Research Institute for Software Technology + Paris Lodron University Salzburg

Chair of the RIST++ ist Prof. Peter Zinterhof. The RIST++ is located at the University of Salzburg and at the Technology Center of Salzburg.

The RIST++ was founded in 1986 as ‘Forschungsinstitut für Softwaretechnologie’. Currently the RIST++ consists of the following labs:

• Parallel Numeric and Supercomputing-Lab

• Basic Research-Lab

• Image-Lab

• Fuzzy-Lab

• Signals and Systems-Lab

• Applied Number Theory-Lab

• Neural Computation-Lab

In each lab problems in basic research and applied research for implementation to sequential and parallel computers are the topics. Some of the labs are partly financed by the industry. The equipment used consists of some transputer systems (now mainly for minor projects), the CONVEX 220 situated at the Technology Center Salzburg, and a Cluster consisting of 7 DEC 3000 Alpha AXP 400 workstations.

3.6 List of Documentation that is available at each ACPC Group

The following sections show for each group the list of documentation that has been copied and sent to all other ACPC groups. So, you can see here the documentation that is available at all ACPC groups. For more detailed documentation you have to contact the network coordinator of the group where the software product is installed.

3.6.1 RISC-Linz

• Guide to Parallel Programming on Sequent Computers

• Programming Guide (ch. 1, 2, 4, index)

• C & Language Tools

– C Users Manual

– pdbx Users Manual

• Reference Manual – About this Manual

– Section 3 (PPS)

– Permuted Index

• μ-System Annotated Reference Manual

• STRAND\underline{88}Release Notes for Users (Sequent Symmetry)

• STRAND\underline{88}Technical Description – Admirality Release (Aug.89)

• C-Linda Reference Manual

3.6.2 Institute for Applied Computer Science and Information Sy-

stems, University of Vienna

• “PARIX Release 1.1. Software Documentation”, Parsytec Computer GmbH, D-5100 Aachen, Juelicher Strasse 338, Germany, 1992.

• “PARIX Release 1.1. Manual Pages”, Parsytec Computer GmbH, D-5100 Aachen, Juelicher Strasse 338, Germany, 1992.

3.6.3 TU-Vienna

• iPSC/860 Users Guide

• iPSC/860 Network Queueing Systems Manual

• iPSC/860 Debugger Manual

• iPSC/860 Programmers Reference Manual

• iPSC/860 Concurrent Programming Vol. 3

• iPSC/860 Simulators Manual

• iPSC/860 Technical Documentation Guide

• iPSC/860 Fortran Compiler Users Guide

• iPSC/860 C Compiler Users Guide

3.6.4 GUP-Linz

• “nCUBE 2 Programmer’s Guide”, nCUBE Corporation, 1990.

• “nCUBE 2 Programmer’s Reference Manual”, nCUBE Corporation, 1990.

• “nCUBE 2 Programmer’s Reference Manual”, nCUBE Corporation, 1990.

• “nCUBE 2 Math Library Manual: Introduction and SSS32”, nCUBE Corporation, 1990.

• “nCUBE 2 Math Library Manual: SSS64 and BLAS”, nCUBE Corporation,

Facilities in the Network

This chapter contains a group description for each of the ACPC groups. This contains a description of the hardware that is available at each of these groups and of the software that is installed at each of these computers. The whole description is divided in several sections, one for each ACPC group. Note that Appendix A contains a table of the machines that are described in more detail in this chapter and that Appendix B contains a list of all persons in the ACPC. These appendices summarize information that is given in a different form in this chapter.

4.1 Group Description of RISC-Linz

General Information about the Group

• Group Name: Parallel Computation Group at the Research Institute for Symbolic Computation (RISC-Linz).

• Group Leader: Prof. B. Buchberger.

• Address:

RISC-Linz

Research Institute for Symbolic Computation Johannes Kepler University A-4040 Linz, Austria, Europe

• Phone Number: 07236-3231 (secretary: 07236-3231-21).

• FAX Number: 07236-3231-30

• Domain Name: Internet: risc.uni-linz.ac.at. Bitnet: ALIJKU11.

• Persons:

Name  Internet-account Bitnet-account Phone Number
Bruno Buchberger  buchberg  K313370  41
Hoon Hong  hhong  K314577  48
Tudor Jebelean  tjebelea  —  64
Hans Wolfgang Loidl  hwloidl  K318576  66
Wolfgang Schreiner  schreine  K315640  66
Kurt Siegl  siegl  K316340  62
Dan Teodosiu  danteo
Peter Zettler  zettler  K312970  67
Werner Landerl  landerl  —  66

4.1.1 Sequent Symmetry S81

Footprint

System  Sequent Symmetry S81
Architecture  Shared Memory MIMD
# of Processors  20
Type of Processors Intel 80386/80387
Memory Size  128 MB
Disk Capacity  4.32 GB
Operating System DYNIX/ptx 2.0

(based on UNIX System V R 3.2)

Name  melmac.risc.uni-linz.ac.at
Internet Address  193.170.36.100
Location  RISC
System Administrator K. Erdei (kerdei)
Network Coordinator H. W. Loidl (hwloidl)

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
good  good  slow good good  good

As the 64Kb/s connection between Linz and Hagenberg is available now, the connection to the machine from outside RISC should be quite good.

Further Description

This machine is not only used for research purpose but also as the main computing server for the institute. Therefore, computational intensive jobs should be run during the night hours or at weekends.

Software

μSystem

• Name: μSystem 4.4.2

• Location: /usr/local/usystem

• Type: Parallel Programming Package.

• Characterization: Library of C routines (and C structures) for handling lightweight tasks. Designed for Shared-Memory Machines. Furthermore: Synchronization via Semaphores and Message Passing. Non-blocking I/O is supported. Exception handling is implemented.

• Functions:

    – Invocation: /usr/local/bin/concc, /usr/local/bin/parcc.

– Termination:

    – Help: man concc, man parcc.

– Example: (Path has been set to /usr/local/bin). Compile program for single-process execution:

      concc /usr/local/usystem/examples/MessagePassing/ProdCons.c

Compile program for multi-process execution:

      parcc /usr/local/usystem/examples/MessagePassing/ProdCons.c

– Other important functions: For monitoring of programs use the following

      compiler: /usr/local/bin/moncc

• Installation:

• Manuals: Manual pages for concc, parcc and moncc. “μSystem Annotated Reference Manual” (also available as ‘.ps’ file).

aXe

• Name: aXe 3.1

• Location: /usr/local/bin/axe (/usr/local/lib/X11/aXe)

• Type: ASCII-Text editor

• Characterization: aXe is a text editor for program source code or LaTeX input or simple text files. It is NOT a WYSIWYG editor. aXe features multiple windows, multiple buffers per window, a file selection box, a so called “minibuffer” which allows to call external commands. It allows cut and paste, and has many functions selectable from menus.

• Functions:

– Invocation: axe

– Termination: There’s a Quit button.

– Help: There’s a help menu with the sub-topics General Help, Key Bindings, Customization, Information about Popups.

– Example:

– Other important functions: It is possible to go to a line with a specific

number.

• Installation: Runs only under X Windows.

• Manuals: man axe, but this mostly refers to the Help menu.

Bash • Name: GNU Bourne Again SHell 1.12

• Location: /usr/local/gnu/bash-1.12

• Type: A bourne shell compatible command language interpreter

• Characterization: Extension of the traditional bourne shell that also includes most of the features of the c-shell and of the korn-shell. Furthermore, this shell allows command line editing with vi-like commands. Implementation of IEEE Posix Shell and Tools specification.

• Functions:

    – Invocation: /usr/local/bin/bash

– Termination: exit

– Help: help

    – Example: bash -rcfile ~/.cshrc

Start the bash and use the file ‘~/.cshrc’ as startup file (Note: As c-shell and bash scripts are not fully compatible some errors will probably occur).

– Other important functions:

• Installation: Put the setting of your environment in the file ‘~/.bashrc’.

• Manuals: man pages; a paper describing the features of bash can be found in ‘/usr/local/gnu/bash-1.12/documentation’.

Bison • Name: GNU Bison 1.14

• Location: /usr/local/gnu/bison-1.14

• Type: Parser generator

• Characterization: yacc replacement

• Functions:

    – Invocation: /usr/local/bin/bison

– Termination:

    – Help: man bison

– Example:

– Other important functions:

• Installation:

• Manuals: Bison Reference Manual (bison.texinfo), man pages

Crisp • Name: Crisp 1.9

• Location: /usr/local/crisp

• Type: Editor

• Characterization: DOS Brief look alike Editor

• Functions:

    – Invocation: /usr/local/bin/cr

– Termination: ALT-x

– Help: ALT-h

– Example:

– Other important functions: Multiple windows, programming language sup-

port, mailing, copy, cut and paste of columns

• Installation: Set environment vars by: source ~crisp/crisp.setenv

• Manuals: Interactive help system DYNIX/ptx

• Name: Dynix/ptx 2.0

• Location: melmac.risc.uni-linz.ac.at

• Type: UNIX System V Release 3.2.0 i386

• Characterization: Standard System V operating system with parallel extensions for a shared memory parallel computer.

• Functions:

    – Invocation: rlogin melmac

– Termination: exit

– Help: help

– Example:

– Other important functions:

• Installation: Ask for an acount by email to schreine.

• Manuals: UNIX System On-Line Help, man pages, DYNIX/ptx reference manual, DYNIX/ptx user guide

Emacs • Name: GNU Emacs 18.59

• Location: /usr/local/gnu/emacs

• Type: Editor

• Characterization: Fully featured editor

• Functions:

    – Invocation: /usr/local/bin/emacs

– Termination: CTRL-x CTRL-c

    – Help: CTRL-h

– Example:

– Other important functions: Multiple windows, programming language sup-

port, mailing

• Installation:

• Manuals: Interactive Help system, man pages Flex

• Name: GNU Flex 2.3 26

• Location: /usr/local/gnu/flex-2.3.7

• Type: Lexical analyzer generator

• Characterization: Tool for generating scanners: programs which recognize lexical patterns in text.

• Functions:

    – Invocation: /usr/local/bin/flex

– Termination:

    – Help: man flex

– Example:

– Other important functions:

• Installation:

• Manuals: flex, flexdoc man pages

GNU Assembler • Name: gas 1.38

• Location: /usr/local/gnu/gas-1.38

• Type: i386 Assembler

• Characterization: Downward compatible to unix “as”.

• Functions:

    – Invocation: /usr/local/bin/a386

– Termination:

– Help: man as

– Example:

– Other important functions:

• Installation:

• Manuals: “as” manual descripton GNU C

• Name: gcc 2.3.3

• Location: /usr/local/gnu/gcc-2.3.3

• Type: C-Compiler

• Characterization: Fast Ansi C Compiler

• Functions:

    – Invocation: /usr/local/bin/gcc

– Termination:

    – Help: man gcc

– Example:

– Other important functions:

• Installation:

• Manuals: gcc man page

GNU Locate • Name: glocate

• Location: /usr/local/bin/glocate

• Type: Fast file find tool

• Characterization: Searches a database for pathnames that match the pattern that has been given on the command line. Currently, the database contains all files on melmac except those in temporary directories like /tmp and /usr/tmp. The database is rebuilt every day at 4 am.

• Functions:

    – Invocation: /usr/local/bin/glocate

– Termination:

    – Help: man glocate
    – Example: glocate *.tar.Z

Lists all compressed tar files.

– Other important functions:

• Installation:

• Manuals:

GNU Zip

• Name: gzip 0.8.1

• Location: /usr/local/gnu/gzip-0.8.1

• Type: File compression and expansion

• Characterization: The GNU Zip tools gzip, gunzip, zcat compress and decompress files by using Lempel-Ziv coding (LZ77). The usage of gzip and gunzip as quite similar to compress and uncompress. Compressed files get the extension .z. This compression method is better than the standard compress. gzip can be used as filter if no file is specified.

• Functions:

    – Invocation: /usr/local/bin/gzip

– Termination:

    – Help: man gzip, /usr/local/bin/gzip -h
    – Example: gzip -v -9 foo

Compresses the file foo yielding a file foo.z. For that the best compression method is used (-9) and the compression factor is printed due to verbose mode (-v).

      gunzip foo.z or gzip -d foo.z

Decompress the file that has been compressed above.

– Other important functions: Multiple compressed files can be concatenated. Note: There are also various other (de-)compression tools available like: zip (DOS port), zoo, squeeze0 and unarj. See the corresponding man pages for further information.

• Installation:

• Manuals:

Gopher Client

• Name: gopher, xgopher version II release 5

• Location: /usr/local/unixutil/gopher+1.2b2

• Type: ASCII and X11 client for accessing gopher servers.

• Characterization: These clients allow the communication with arbitrary gopher servers. The Internet Gopher is a distributed document delivery service. The documents are organized in hierarchical way. The Internet Gopher Servers accept simple requests from the clients and respond by sending the client a document.

• Functions:

    – Invocation: /usr/local/bin/gopher, /usr/local/bin/xgopher

– Termination: Quit Button

    – Help: Help Button, man gopher, man xgopher
    – Example: xgopher melmac.risc.uni-linz.ac.at

This will connect the user with the gopher server at RISC-Linz. The user can then walk through the hierarchy of available files in the same way as in a usual directory hierarchy. Documents can also be directly fetched from the server.

– Other important functions:

• Installation: For gopher: setenv PAGER /usr/local/bin/less. Several other

environment variables might be set (e.g. for defining with which program mail should be sent). See the man page of gopher. One can create an own ~/.gopherrc file containing bookmarks, configuration information etc.

• Manuals:

Less Paginator • Name: Less

• Location: /usr/local/less

• Type: File displaying tool.

• Characterization: Similar to “more”, but allows backwards movement.

• Functions:

    – Invocation: /usr/local/bin/less

– Termination: q

– Help: h

– Example:

– Other important functions:

• Installation:

• Manuals: less man page, online help Maple

• Name: MAPLE V

• Location: /usr/local/maple

• Type: Interactive Computer Algebra System

• Characterization: Maple has the ability to algebraically manipulate unbounded integers, exact rational numbers, real numbers with arbitrary precision, symbolic formulae, polynomials, sets, lists, and equa- tions. It can solve systems of equations and differentiate and integrate formulae. Note: There is also a Maple version with an X11 interface: xmaple.

• Functions:

    – Invocation: /usr/local/bin/maple

– Termination: quit

– Help: ?

– Example:

– Other important functions:

• Installation:

• Manuals: Maple Reference Manual, Interactive Help, man pages

∥MAPLE∥ • Name: ∥MAPLE∥ Shell 1.0

• Location: /usr/local/maple/pmaple

• Type: Interactive shell of ∥MAPLE∥ (Parallel Maple), Parallel Computer Algebra System

• Characterization: Computer Algebra System, in which logic programming provides parallelism and imperative programming provides efficiency. The system is built as a combination between the parallel declarative programming language Strand and the sequential computer algebra system Maple. The system supports the whole Maple library and all Strand features.

• Functions:

    – Invocation: /usr/local/bin/parmaple

– Termination: <CTRL> C

    – Help: man parmaple

– Example:

– Other important functions:

• Installation:

• Manuals: man pages, /usr/local/maple/man/pmapledoc.dvi - the documentation

Netrek • Name: XtrekII 6.1

• Location: /usr/local/netrek

• Type: Game

• Characterization: Multi-users X-windows Star Trek game

• Functions:

    – Invocation: /usr/local/bin/netrek

– Termination: Q

– Help: h

– Example:

– Other important functions:

• Installation: A private environment can be set by creating an own ~/.netrekrc file.

• Manuals: Online start-up instructions

PACLIB • Name: PACLIB (Version 1.18)

• Location: /usr/local/paclib

• Type: Parallel System for Computer Algebra.

• Characterization: PACLIB consists of a compiler for a parallel programming language (based on C) and a library of sequential and parallel computer algebra algorithms (based on SACLIB). The user can implement parallel algorithms and monitor their dynamic behavior using the visualization tools pacgraph and

  pacutil.

• Functions: The PACLIB compiler is invoked as pacc (no optimization) or paco (optimization) or pacd (debugging) with the same syntax as the GNU C compiler. The final program can be called with several runtime options e.g.

    pacc test.c -o prog
    prog +p16 +N1000000 +t1

The option +p defines the number of processors, +N defines the number of heap words, and +t1 triggers the generation of profiling information that is written into a file pac.dat. This information can be analyzed by the tools pacgraph and pacutil. The commands

    pacgraph 250 250 <pac.dat >grap.tex
    pacutil  250 250 <pac.dat >util.tex

generate two LaTEX pictures of size 250 times 250 point that display the activity of the tasks and the utilization of the processors, respectively.

• Manuals: PACLIB User Manual (RISC-Linz Technical Report 92-32 and ACPC Technical Report ACPC/TR 92-9).

Patch • Name: GNU Patch 2.0.12u4

• Location: /usr/local/gnu/patch-2.0.12u4

• Type: Program patching tool

• Characterization: Applies a diff file to an original.

• Functions:

– Invocation: /usr/local/bin/patch orig patchfile

– Termination:

    – Help: man patch

– Example:

– Other important functions:

• Installation:

• Manuals: patch man page

Perl • Name: GNU Perl 4.019

• Location: /usr/local/gnu/perl-4.019

• Type: Practical Extraction and Report Language • Characterization: An interpreted language optimized for text processing like

scanning text files and extracting information. It combines features of C, sed, awk and sh.

• Functions:

    – Invocation: /usr/local/bin/perl

– Termination: <CTRL> C

    – Help: man perl

– Example:

– Other important functions:

• Installation:

• Manuals: Perl Reference Guide, Frequently Asked Questions about Perl (both in ‘/usr/local/gnu/perl-4.019’).

Raytracer • Name: Rayshade V3

• Location: /usr/local/ray3

• Type: Ray tracing program

• Characterization: Rayshade reads a file describing a scene to be rendered and produces a Utah Raster RLE format file of the ray-traced image.

• Functions:

    – Invocation: /usr/local/bin/rayshade

– Termination:

    – Help: man rayshade

– Example:

– Other important functions:

• Installation:

• Manuals: rayshade man page

SACLIB • Name: SACLIB 1.1, ISAC 2.0

• Location:

– /usr/local/saclib: The maintenance directory.

– /usr/local/saclib/saclib1.1: The root of the directory tree for the current version.

• Type: Computer Algebra Library for the C programming language.

• Characterization: SACLIB contains programs for list processing, infinite precision arithmetic (integer, rational and modular), operations on multivariate polynomials (arithmetic operations, substitution, greatest common divisors, resultants, factorization, and others), polynomial real root isolation and refinement, and operations with real algebraic numbers and polynomials having algebraic number coefficients.

There is also a simple interactive shell called isac which can be used for experimenting with the functions in the library.

• Functions:

These are only the functions of isac, the interactive shell. In general, the library functions are intended to be called from a program.

– Invocation: isac

– Termination: quit;

– Help: help;

– Example:

        1> I := 50;
        2>  F := IFACTL(I);
        3> IWRITE(F);
        30414093201713378043612608166064768844377641568960512000000000000
        4> quit;

– Other important functions: The shell scripts sman and sdesc can be used for getting help on SACLIB functions. See the manual for a description of these commands.

• Installation: Your environment must have the variable saclib set to

/usr/local/saclib/saclib1.1 and your path must contain
/usr/local/saclib/saclib1.1/bin.
• Manuals: /usr/local/saclib/saclib1.1/doc/saclib.dvi and
/usr/local/saclib/saclib1.1/doc/saclocal.dvi.

STRAND\underline{88}

• Name: STRAND\underline{88} Buckingham Release • Location: /usr/local/strand

• Type: Parallel Declarative Programming Language

• Characterization: STRAND\underline{88} is an implementation of the Strand language, which includes an environment for developing and running Strand programs. The Strand language is desinged to utilize the facilities of concurrent machines and parallel architectures. It supports concurrent processing, stream communication between processes, Data-flow, modular development, declarative reading, object oriented programming and interfaces to other languages and data types.

• Functions:

– Invocation: st

– Termination: <CTRL>C

– Help: st -h

– Example:

– Other important functions:

• Installation: source /usr/local/strand/strandrc

• Manuals: STRAND\underline{88} User Manual, Book: Strand New Concepts in Parallel Programming

TeX, LaTeX • Name: TeX C Version 3.14t3

• Location: /usr/local/tex

• Type: Text formatting and typesetting system

• Characterization: Standard Scientific Word processor

• Functions:

    – Invocation: /usr/local/bin/tex, /usr/local/bin/latex

– Termination:

    – Help: man tex, man latex

– Example:

– Other important functions:

• Installation: Set environment vars by: source ~tex/lib/TeX3.setenv

• Manuals: LaTex - A Document Preparation System by Leslie Lamport The TeXbook by Donald E. Knuth, man pages TEXCAD

• Name: TeXcad 1.2

• Location: /usr/local/tex/src/xtexcad

• Type: Simple interactive tool for the generation of LaTEX pictures.

• Characterization: The program allows you to create in an interactive way LaTEX pictures consisting of straight lines, arrows, boxes, circles and text. Copy and erase operations are supported.

• Functions:

– Invocation: texcad

– Termination: Click on button FILE and menu item TERMINATE.

    – Example: texcad test.pic
• Manuals: /usr/local/tex/src/xtexcad/texcad12.man (LaTEX file).

Tgif • Name: Tgif Version 2.12-p7

• Location: /usr/local/lib/tgif

• Type: X-lib based 2-D drawing facility under X11.

• Characterization: Tgif is an interactive drawing tool that allows the user to draw and manipulate objects in the X Windows System (it is based on Xlib). Tgif supports a bottom-up construction of hierarchical objects as well as a topdown specification of drawings. Among the formats that can be incorporated into a drawing are X11 bitmaps, Encapsulated Postscript and many others. The supported output formats are Postscript, Encapsulated Postscript, X11 bitmap and X11 pixmap.

• Functions:

    – Invocation: /usr/local/bin/tgif

– Termination: <CTRL> Q

    – Help: man tgif
    – Example: tgif foo.ps

Allows the user to modify the postscript file foo.ps.

– Other important functions:

• Installation: The following environment variables should be set:

    – setenv TGIFPATH /usr/local/lib/tgif,
    – setenv TGIFICON tgificon.obj.

• Manuals:

Utah Raster Toolkit • Name: Utah Raster Toolkit 3.0

• Location: /usr/local/urt

• Type: C graphic library

• Characterization: The Utah Raster Toolkit is a collection of programs and C routines for dealing with raster images commonly encountered in computer graphics. A device and system independent image format stores images and information about them.

• Functions:

– Invocation: include functions to your C-programs

– Termination:

    – Help: man urt

– Example:

– Other important functions:

• Installation:

• Manuals: urt man pages

X-Windows • Name: ptx/windows X11 R4, R5

• Location: /usr/bin/X11, /usr/local/X11R5

• Type: Portable, network-distributed window system

• Characterization: A window system developed at MIT. X uses a client-server architecture to promote portable applications.

• Functions:

– Invocation:

– Termination:

– Help: xman

– Example:

– Other important functions: OSF-Motiv Window manager (mwm)

• Installation:

• Manuals: ptx/windows Programmers Reference, ptx/windows Users Reference, man pages, The X Windows System Xmandel

• Name: xmandel, pxmandel

• Location: /usr/local/x11/xmandel

• Type: Demonstration program.

• Characterization: X11 interface for generating and colouring Mandelbrot and Julia sets. There is a sequential program (xmandel) as well as a parallel program (pxmandel) available, that can be used for demonstrating the speed-up achievable by parallel computation.

• Functions:

    – Invocation: xmandel, pxmandel <no. of processors>

– Termination: Quit Button

    – Help: man xmandel
    – Example: pxmandel 16 -geometry 400x450+450+0 &

Creates a window that is a little smaller than half of the monitor size. The parallel computation is started by clicking at the ‘Mandelbrot’ or ‘Julia’ button.

       pardemo/pxdemo

Starts two windows: One for sequential computation and one for parallel computation on 16 processors. This shell script is highly recommended for demos! Note: Don’t move the cursor inside the pictures while they are computed!

– Other important functions: By clicking at the zoom button and then at a part of the picture, this part will be zoomed and recomputed.

• Installation: Don’t forget to set your DISPLAY environment variable before starting the program.

• Manuals:

XParty

• Name: XParty (XPhone)

• Location: /usr/local/x11/xparty

• Type: Communication Tool

• Characterization: Multi-user/multi-display communication utility. Allows several people to perform a “Network Conference”. The user who starts the conference leads the discussion and has several privileges. Motif based. Can also be used for one-way communication.

• Functions:

– Invocation: /usr/local/bin/xparty

– Termination: Quit-Button

– Help: help-Button

– Example: Start program with xparty. Press AddR-Button to add a recipient. Select the new recipient by pressing the corresponding button in the ‘Recipients’ part of the screen. If the recipient accepts the call he will be added to the ‘on the air’ list and the communication can begin.

– Other important functions: One way communication by pressing the one-way button before sending.

• Installation: Parameters can be set in your own ‘~/.Xdefaults’ file. See the file‘/usr/lib/x11/app-defaults/XParty’ for the default settings. The hidden file ‘~/.xmsg hosts’ contains a list of recipients (in the format ‘label@display’). When a new recipient is added in xparty, a new entry is automatically created in this file.

• Manuals: xparty man page

4.1.2 Transputer Network

Footprint

System  Impuls Transputer Network
Architecture  Distributed Memory MIMD
# of Processors  16
Type of Processors T800
Memory Size  16 * 4 MB = 64 MB
Disk Capacity  0 MB
Operating System Helios (optionally !)
Name  —
Internet Address  —
Location  RISC
System Administrator K. Siegl (siegl)
Network Coordinator H. W. Loidl (hwloidl)

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
good  good  slow good good  good

Further Description

The transputer network is not always turned on. Contact the network administrator if you want to work on it. This transputer network has no fixed topology. Before starting the actual application the user has to set up a special topology.

Footprint of the Host Computer

System  AT-PC
Type of Processors Intel 80386
Memory Size  8 MB
Disk Capacity  100 MB
Operating System Interactive System V R 3.2
Name  venus.risc.uni-linz.ac.at
Internet Address  193.170.36.70
Location  RISC
System Administrator W. Schreiner (schreine)
Network Coordinator H. W. Loidl (hwloidl)

Software

ANSI C Toolset

• Name: ANSI C Toolset of INMOS (Version IMS D711)

• Location: /usr0/dos/ictools/tools

• Type: C compiler, linker and configurer

• Characterization: Like the occam toolset this toolset contains several programs for programming in ANSI C on a transputer. Among others, this toolset contains the following programs:

– icc — the ANSI C compiler,

– ilink — the toolset linker,

– icconf — the configurer tool,

– icollect — the code collector.

• Functions: For the most important tools listed above, their usage is described in more detail here:

– icc — the ANSI C compiler. An ANSI C Compiler with concurrency support of INMOS that has been developed from the 3L Parallel C Compiler, which implemented K & R C.

∗ Invocation: icc filename { options } ∗ Termination: <CTRL> C

∗ Help:

∗ Example: icc foo.c Compiles the C program in foo.c and creates the result in foo.tco.

∗ Other important functions:

– ilink — the toolset linker. The toolset linker for linking separately compiled program units and libraries into a single code unit.

∗ Invocation: ilink { filenames } { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: ilink foo1.tco foo2.tco

Links the compiled C programs foo1.tco and foo2.tco into the file

  foo1.lku.

∗ Other important functions:

– icconf — the configurer tool. The configurer tool generates configuration information for transputer networks from a configuration description written in the transputer configuration language.

∗ Invocation: icconf filename { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: icconf foo.cfs

Produce the configuration data file foo.cfb from the configuration description file foo.cfs. The foo.cfb file can then be taken as the input

  file for icollect.

∗ Other important functions:

– icollect — the code collector. The code collector generates executable files for single and multitransputer programs, from configuration data files and linked units respectively.

∗ Invocation: icollect filename { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: icollect foo.lku /t

This creates a bootable file for a single tranpsuter in foo.btl.

  icollect foo.cfb

(This creates a file foo.btl that can be executed on a multitransputer network).

∗ Other important functions:

• Installation: SET IBOARDSIZE=#400000 (this sets the transputer memory size for the bootstrap loader).

SET TRANSPUTER=#150 (this sets the transputer base address).

  SET ITERM=Z:\USR0\DOS\ITOOLS\ITERMS\IBMPC.ITM (the terminal description

file for a PC).

  SET ISEARCH=Z:\USR0\DOS\ITOOLS\LIBS (this sets the library path).

• Manuals: “ANSI C Toolset — User Manual”, by INMOS Limited, 1990.

Helios • Name: Helios 1.2.1

• Location: /usr0/dos/helios

• Type: Parallel Operating System

• Characterization: Helios is an Unix look alike Operating system for distributed memory machines (i.e. Transputers)

• Functions:

– Invocation: /usr0/dos/helios/helios0 - Single transputer /usr0/dos/helios/hyper - (Hypercube) Network version.

    – Termination: <CTRL><SHIFT>F9 or

Login under user: shutdown

– Help: help

– Example:

– Other important functions:

• Installation: Start up VPIX (DOS emulation) first.

  SET IBOARDSIZE=#400000

• Manuals: Helios User Manual, man pages

LSC Transputer Toolset • Name: Logical Systems Transputer Toolset (Version 89.1.c)

• Location: /usr0/lsc

• Type: C preprocessor, compiler, assembler, linker.

• Characterization: Like the ANSI C toolset this toolset contains several programs for programming in C on a transputer. Among others, this toolset contains the following programs:

– pp — the C preprocessor,

– tcx — the C compiler, – tasm — the assembler,

– tlnk — the linker,

– tlib — the relocatable file librarian,

– make — a program maintenance utility,

– ld-one — the single transputer loader,

– ld-net — the network loader for arrays of transputers,

– cio — the host I/O driver for C I/O on the transputer,

– tio — a simple example driver for standard I/O.

• Functions: For the most important tools listed above, their usage is described

in more detail here:

– pp — the C preprocessor. This C preprocessor conforms to the ANSI Draft Proposed “C” standard and contains options for backward compatibility with programs written for non-ANSI preprocessors.

∗ Invocation: pp filename { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: pp foo.c -o foo

The preprocessor takes foo.c as input and creates the file foo as output. The default suffix for the produced file is ‘.pp’.

∗ Other important functions:

– tcx — the C compiler. Cross compiler for the INMOS 16 and 32 bit Trans-

puters. It does not include preprocessing.

∗ Invocation: tcx filename { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: tcx foo -v -p8

Compile the program foo.pp in verbose mode (option -v) for a T800 Transputer (option -p8). The created file is foo.tal.

∗ Other important functions:

– tasm — the transputer assembler. A relocating assembler for INMOS Transputers. It supports standard INMOS mnemonics and allows splitting a program into separate pieces, which are combined at linkage time.

∗ Invocation: tasm filename [temp-directory] { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: tasm foo /fast

Assemble the program foo.tal and use the directory fast for storing the temporary files. The created file is foo.trl.

∗ Other important functions:

– tlnk — the transputer linker. This linker has been designed for use with INMOS Transputers. It supports multiple libraries and allows complete control of load address for each program or library code fragment.

∗ Invocation: tlnk [ command-file ]

∗ Termination: <CTRL> C

∗ Help:

∗ Example: tlnk By this call the linker will ask for the files to be linked, for some flags, load address, stack address and other parameters.

∗ Other important functions:

– ld-one — the single transputer loader. This tool is used to load an executable file onto a single transputer. Executable files are created as a result of linking using tlnk (default suffix ‘.tld’). The first argument is the file to be downloaded. The second file runs on the host and handles transputer I/O requests.

∗ Invocation: ld-one filename filename

∗ Termination: <CTRL> C

∗ Help:

∗ Example: ld-one foo tio

This starts foo.tld on a single transputer and tio on the host for I/O handling.

∗ Other important functions:

– ld-net — the network transputer loader. This tool is used to load a collection of executable files onto a network of transputers. Executable files are created as a result of linking using tlnk (default suffix .tld). The first argument is the file specifying the file to read to determine network topology and the list of programs to be loaded. The remaining parameters on the command line are passed to the host driver program.

∗ Invocation: ld-net network-info-file arguments

∗ Termination: <CTRL> C

∗ Help:

∗ Example: ld-net load

Take the network-information-file load.nif and load the program, that is specified in load, on a network, whose topology is also specified in load. The program will be executed with the program cio.exe handling the I/O.

∗ Other important functions:

• Installation: SET TMP=Z:\USR\TMP (directory for temporary files).

SET PPINC=Z:\USR0\LSC\TINCLUDE (directory for include files).
SET TLIB=Z:\USR0\LSC\TLIBRARY (directory for libraries).

Add the following directory to your path: Z:\USR0\LSC\DOS.

• Manuals: “Logical Systems Transputer Toolset”, by Logical Systems, 1989.

The Network Configurer • Name: Transputer Network Configuration Program.

• Location: /usr0/dos/tutil

• Type: Utility for fixing a given topology.

• Characterization: This is an occam program (so it must be started with the iserver) for fixing the topology of the transputer network that is described in the topology file with which this program is called. Topology descriptions for a network with up to 16 transputers can be found in /usr0/dos/tutil/topo.

• Functions:

    – Invocation: /usr0/dos/itools/tools/iserver /sb
      /usr0/dos/tutil/conf.b8h topology

– Termination: <CTRL> C

– Help:

    – Example: iserver /sb /usr0/dos/tutil/conf.b8h
      /usr0/dos/tutil/topo/torus.top

This fixes the topology of a torus of 16 transputers. /usr0/dos/itools/tools must be in the path.

– Other important functions:

• Installation: /usr0/dos/itools/tools should be put into the path.

• Manuals:

Occam 2 Toolset • Name: Occam 2 Toolset for IBM/NEC PC by Inmos (Version D705B).

• Location: /usr0/dos/itools/tools (of host PC)

• Type: A collection of independent programs for programming in occam.

• Characterization: This toolset consists of various programs that allow the user to work with a transputer (network). The toolset contains the following programs:

– iboot — the bootstrap tool,

– icheck — the occam 2 syntax checker,

– iconf — the configurer, – idebug — the debugger,

– idump — the memory dumper,

– ilibr — the librarian,

– ilink — the linker,

– ilist — the binary lister,

– imakef — the makefile generator,

– iserver — the host file server,

– isim — the T414 simulator,

– iskip — the skip loader tool,

– occam — the occam 2 compiler.

• Functions:

For the most important tools listed above, their usage is described in more detail here:

– iboot — the bootstrap tool. This tool produces executable code for single transputers. The object file it generates can be loaded onto a single transputer and run. The input must be a compiled (.txx) or a linked (.cxx) object file.

∗ Invocation: iboot filename { options }

∗ Termination: <CTRL> C

∗ Help:

    ∗ Example: iboot foo.c8h

This takes the compiled and linked file foo.c8h and creates the file foo.b8h that can be started with the iserver on one transputer.

∗ Other important functions:

– icheck — the occam 2 syntax checker. This tool performs a syntax check of the full occam 2 product language, as described in the “Occam 2 Refernece Manual”, but produces no object code. It generates more data and displays more information about the errors than the compiler.

∗ Invocation: icheck filename { options }

∗ Termination: <CTRL> C

∗ Help:

    ∗ Example: icheck foo.occ

This checks the syntax of the occam program in foo.occ. The ‘.occ’ suffix may be omitted.

∗ Other important functions:

– iconf — the configurer. This tool allocates occam processes to processors and assigns channels to links on transputer networks. It takes a configuration description and produces either an object code file ready for booting into a network of transputers, or a configuration map describing the allocation of code and placement of channels.

∗ Invocation: iconf filename { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: iconf foo.pgm

This configures a transputer network that is described in the file foo.pgm. The ‘.pgm’ suffix may be omitted.

∗ Other important functions:

– idebug — the debugger. This tool allows to debug an occam program even when it is executed on a transputer network. With this debugger errors can be located in the source code, parameters can be examined channels and processes can be analysed and procedures can be traced.

∗ Invocation: idebug bootablefile { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: idebug foo.btl

This starts the debugger on the program foo.btl that runs on a transputer network. Possible suffices are ‘.bxx’, ‘.btl’ and ‘.btr’.

∗ Other important functions:

– idump — the memory dumper. This tool dumps the contents of the root processor’s memory to disk. It is used to enable the debugging of code running on the root transputer.

∗ Invocation: idump filename memorysize { startoffset length }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: idump foo.btl #200000 #300000 #100000

This dumps the first half of the memory of the root transputer (bytes 0 – #200000). Additionally, #100000 bytes starting from address #300000 are dumped.

∗ Other important functions:

– ilink — the linker. This tool creates single object files from separately

compiled code and libraries (all external references are resolved). ∗ Invocation: ilink { filenames } { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: ilink foo1.t8h foo2.t8h string.lib

This links the compiled occam files foo1.t8h and foo2.t8h with the routines from the string library in string.lib. The resulting file has

  the name foo1.c8h.

∗ Other important functions:

– iserver — the host file server. This tool loads programs onto transputers and transputer networks and provides the run-time environment through which programs communicate with the host.

∗ Invocation: iserver { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: iserver /sb foo.btl

This loads the program foo.btl on a tranputer network and boots the program (/sb option). To get a file like foo.btl the source files have to be compiled and linked and made bootable by using the bootstrap tool iboot (for single transputer programs) or the configurer iconf (for multitransputer programs). Furthermore, before the iserver can be started the topology of the network has to be fixed with a call to the network configurer (see below).

∗ Other important functions:

– occam — the occam 2 compiler. This tool compiles occam source code programs into the assembler code of the transputer (optionally to T212, T222, M212, T414, T425, T800).

∗ Invocation: occam filename { options }

∗ Termination: <CTRL> C

∗ Help:

∗ Example: occam /t8 /h foo.occ

This compiles the occam program foo.occ into the assembler code of the T800 transputer in halt mode (i.e. the transputer halts when an error occurs in the program). The resulting file is foo.t8h.

∗ Other important functions:

• Installation: SET IBOARDSIZE=#400000 (this sets the transputer memory size for the bootstrap loader).

SET TRANSPUTER=#150 (this sets the transputer base address).

SET ITERM=Z:\USR0\DOS\ITOOLS\ITERMS\IBMPC.ITM (the terminal description

file for a PC).

SET ISEARCH=Z:\USR0\DOS\ITOOLS\LIBS (this sets the library path).

• Manuals: “Occam 2 Toolset, User Manual, IBM/NEC PC”, by Inmos Limited, 1989.

“Occam 2 Reference Manual”, by Inmos Limited, Prentice Hall, 1988.

ParaScope Debugger

• Name: ParaScope Transputer Network Debugger for Logical Systems C.

• Location: /usr0/lsc/dos/cdbug

• Type: Debugger for parallel programs.

• Characterization: This debugger works with Logical Systems C on a network of Transputers. It is a ‘postmortem’ debugger. In several windows the source code, the memory contents and other information is shown.

• Functions:

– Invocation: cdbug { options }

– Termination: <CTRL> C for terminating the program that has to be debugged and to enter the debugger with the above command. Go with the mouse to the Process Register Window. Press center mouse button and choose “exit” from the menu that pops up.

– Help:

– Example: cdbug Debug a program that has just crashed on the transputer (network) or that has been stopped with <CTRL> C. In the program that has to be debugged the function debug() has to be called at the beginning of main and for each that will be linked into the main program a symbol table must be generated by calling the program symtab.exe.

– Other important functions: Pressing the center mouse button in different windows causes different menus to pop up.

• Installation: SET CDPATH=Z:\USR0\LSC\DOS\CDBUG (directory for the debugger).

• Manuals: “ParaScope Transputer Network Debugger for Logical Systems C”, by Computer Systems Architects (CSA), 1990.

STRAND\underline{88} for Helios

• Name: STRAND\underline{88} Buckingham Release

• Location: /usr0/dos/strand

• Type: Parallel Declarative Programming Language

• Characterization: STRAND\underline{88} is an implementation of the Strand language, which includes an environment for developing and running Strand programs. The Strand language is designed to utilize the facilities of concurrent machines and parallel architectures. It supports concurrent processing, stream communication between processes, Data-flow, modular development, declarative reading, object oriented programming and interfaces to other languages and data types.

• Functions:

– Invocation: st

– Termination: <CTRL>C – Help: st -h

– Example:

– Other important functions:

• Installation: Start up the Helios environment first.

     source /helios/users/strand/strandrc

• Manuals: STRAND\underline{88} User Manual, Book: Strand New Concepts in Parallel Programming

4.1.3 Silicon Graphics Iris VGX Graphics Supercomputer

Footprint

System  SGI Iris Power Series
Architecture  Shared Memory MIMD
# of Processors  4
Type of Processors MIPS R3000
Memory Size  64 MB
Disk Capacity  2 GB
Operating System IRIX System V.3
Name  edelweiss.risc.uni-linz.ac.at
Internet Address  193.170.33.151
Location  RISC
System Administrator K. Erdei (kerdei)
Network Coordinator H. W. Loidl (hwloidl)

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
good  —  —  —  —  —

Further Description

The installation of this machine has not been completed, yet. Therefore, there are currently no ordinary user accounts on the machine and there are also no guest accounts for the other ACPC groups installed. As soon as this machine will be opened for general usage, the guest accounts will be installed. This will be announced over the ACPC mailing list.

Software

4.2 Group Description of the Institute for Applied Computer Science and Information Systems, University of Vienna

General Information about the Group

• Group Name: ACPC Group at the Institute for Applied Computer Science and Information Systems

• Group Leader: Prof. Dr. Günter Haring

• Address:

University of Vienna

Inst. for Applied Computer Science and Information Systems Lenaugasse 2/8 A-1080 Vienna, Austria, Europe

• Phone Number: ++431 4086366...-10

• FAX Number: ++431 4080450

• Domain Name: Internet: ani.univie.ac.at. Bitnet: awiuni11.bitnet.

• Persons:

Name  Internet-account Bitnet-account Phone Number
Günter Haring  haring  a4424dad  +12
Alois Ferscha  ferscha  a4424dab  +18
Gabriele Kotsis  kotsis  a4424dau  +14
Georg Pigel  pigel  a4424gak  +13
Harald Wabnig  wabnig  a4424gal  +17
Ursula Presslauer  presslauer  —  +23

4.2.1 MultiCluster-3DE

Footprint

System  MC-3DE (Local Development Station)
Architecture  Distributed Memory MIMD
# of Processors  16
Type of Processors T800
Memory Size  16×4 MB = 64 MB
Disk Capacity  0 MB
Operating System Parix Release 1.1
Name  —
Internet Address  —
Location  Lenaugasse 2/8 1080 Vienna
System Administrator H. Wabnig (wabnig)
Network Coordinator H. Wabnig (wabnig)

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
slow  good  good  slow  good

Further Description

Our transputer system is a local development station. Although the parix concept contains software for multiuser administration it is not applicable to the MC-3DE hardware. Therefore these software parts are not included and multiuser processing is not possible. Nevertheless the system can be used over the network. But the system administrator has to be contacted to set up the system accordingly and prevent problems caused by multiple users.

The MC-3DE has a fixed topology (mesh). Special topology configurations can be built with cables.

System  SUN IPX
Type of Processors Sparc
Memory Size  32 MByte
Disk Capacity  1.2 GByte
Operating System SunOS Release 4.1.1-IPX
Name  epikur.ani.univie.ac.at
Internet Address  131.130.32.113
Location  Lenaugasse 2/8 1080 Vienna
System Administrator U. Presslauer (ursula)
Network Coordinator U. Presslauer (ursula)

Software

PARIX

• Name: PARIX Release 1.1

• Location: /home/parix

• Type: Operating System, C-Compiler

• Characterization: PARIX is a new operating system for transputer based multiprocessor systems. It consists of the basic software to load and run programs, software for administration (not included) and ACE EXPERT compilers.

• Functions:

– Example:

Change to the directory: /home/parix/examples. Copy the file hello.c to your home directory. To compile and run the hello.c example type:

ancc.px -o foo.px foo.c
run -f0 1 1 foo.px

There are other examples in the examples directory. For each example exists a information file “doc” which describes how to compile and run the example program.

– Termination: <CTRL> C

– Help: Manual pages are available for all PARIX commands.

• Installation:

For each account a .cshrc file is provided which contains all commands that are necessary to set up the environment.

• Manuals:

“PARIX Release 1.1. Software Documentation”, Parsytec Computer GmbH, D-5100 Aachen, Juelicher Strasse 338, Germany, 1992.

“PARIX Release 1.1. Manual Pages”, Parsytec Computer GmbH, D-5100 Aachen, Juelicher Strasse 338, Germany, 1992.

4.3 Group Description of TU-Vienna

General Information about the Group

• Group Name: ACPC Group at the Technical University of Vienna

• Group Leaders: Dr. W. Kleinert and Univ. Doz. Dr. C. Überhuber

• Address:

Computing Center

Technical University of Vienna Wiedner Hauptstrasse 8-10 A-1040 Vienna, Austria, Europe

• Phone Number: ++431-58801 (secretary: ++431-58801-5481)

• FAX Number: ++431-58-74-211

• Domain Name: Internet: edvz.tuwien.ac.at. Bitnet: awituw64.bitnet

• Persons:

Name  Internet-account Bitnet-account Phone Number
Wolfgang Kleinert  kleinert  kleinert  5480
Irmgard Husinsky  husinsky  husinsky  5484
Josef Fritscher  fritscher  fritscher  5505
Christoph Überhuber  –  –  5412
Arnold Krommer  –  –  –

4.3.1 Sequent Balance B21

Footprint

System  Sequent Balance B21
Architecture  Shared Memory MIMD
# of Processors  28
Type of Processors NS32032
Memory Size  24 MB
Disk Capacity  540 MB
Operating System DYNIX 3.0.14
Name  eacpc1.tuwien.ac.at
Internet Address  128.130.36.3
Location  TU-Vienna
System Administrator J. Fritscher
Network Coordinator J. Fritscher

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
good  good  good

Further Description This machine is used for research and educational purpose.

Software KAP

• Name: Kuck & Assoc KAP/Sequent 7.2

• Location: /usr/bin/kap

• Type: Parallelizing FORTRAN compiler

• Characterization: Fully automatically parallelizing FORTRAN compiler

• Functions:

– Invocation: kap or fortran -K

– Termination:

    – Help: man kap or man fortran
    – Example: fortran -K program.f

– Other important functions: KAP converts IF .. GOTO loops into DO ..

      loops

• Installation:

• Manuals: KAP/Sequent User’s Guide

C-Linda • Name: SCA C-Linda 2.4

• Location: /usr/local/src/linda/v2.4

• Type: Linda pre-compiler with TupleScope debugger

• Characterization: A C compiler using the Linda parallel programming paradigm

• Functions:

– Invocation: clc

– Termination:

– Help: man clc
– Example: clc -linda ts 2000 sample.cl

– Other important functions:

• Installation:

• Manuals: C-Linda Reference Manual

GNU C • Name: gcc 2.3.3

• Location: /usr/local/bin/gcc

• Type: C-Compiler

• Characterization: Fast Ansi C Compiler

• Functions:

– Invocation: gcc

– Termination:

– Help: man gcc

– Example:

– Other important functions:

• Installation:

• Manuals: gcc man page

4.3.2 Intel Hypercube iPSC/860

Footprint

System  Intel Hypercube iPSC/860
Architecture  Distributed Memory MIMD
# of Processors  16
Type of Processors Intel i860
Memory Size  72 MB
Disk Capacity  1.55 GB
Operating System NX Rel 3.3.1
Name  eacpc3.tuwien.ac.at
Internet Address  128.130.36.5
Location  TU-Vienna
System Administrator J. Fritscher
Network Coordinator J. Fritscher

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
good good good

Further Description Footprint of the Host Computer

System  Intel PC 386
Type of Processors i386
Memory Size  8 MB
Disk Capacity  380 MB
Operating System AT&T Unix System V Rel 3.2
Name  eacpc2.tuwien.ac.at
Internet Address  128.130.36.4
Location  TU-Vienna
System Administrator J. Fritscher
Network Coordinator J. Fritscher

Software C-Linda

• Name: SCA iPSC/Hypercube Linda 2.4 • Location: /usr/local/linda

• Type: Linda pre-compiler with TupleScope debugger

• Characterization: A C compiler using the Linda parallel programming paradigm

• Functions:

– Invocation: clc

– Termination:

    – Help: man clc
    – Example: clc -o sample -linda ts 2000 sample.cl
      host -t 4 sample

– Other important functions: A Linda code development system (CDS) running on Sun SparcStations is included with the package. CDS is an Linda system emulator and can be used to develop Linda programs without having a parallel computer.

• Installation: It is necessary to include /usr/local/linda/bin to the path environment variable.

• Manuals: C-Linda Reference Manual

AT&T C++ • Name: CC 2.1

• Location: /usr/local/bin

• Type: C++ Compiler

• Characterization: AT&T’s original C++ Compiler

• Functions:

– Invocation: CC

– Termination:

– Help:

– Example:

– Other important functions:

• Installation:

• Manuals:

PICL (Portable Intrumented Communication Library)

• Name: PICL Version 1/1/92 from netlib

• Location: /usr/local/lib

• Type: Library for parallel programming

• Characterization: In combination with PARAGRAPH an programming and analysis tool is provided. PARAGRAPH can run on any other workstation with an X-Window environment. The libraries can be called from FORTRAN or C.

• Functions:

    – Example: cc -o hostprogram hostprogram.c hostlib.a -host
      icc -o nodeprogram nodeprogram.c nodelib.a -node

• Installation:

• Manuals:

4.4 Group Description of the Theoretical Chemistry Group

General Information about the Group

• Group Name: Theoretical Biochemistry Group

• Group Leader: Prof. O. Schuster

• Address:

Institute for Theoretical Chemistry University of Vienna Währingerstraße 17 A-1090 Vienna

• Phone Number: +43 222 43 61 41 ext. 78

• FAX Number: +43 222 40 28 525

• Domain Name: itc.univie.ac.at, AWIUNI11.BITNET

• Persons:

Name Internet-account Bitnet-account Phone Number
P. Schuster pks@tiger.itc.univie.ac.at A8441DAM 43 61 41 ext. 78
I. Hofacker ivo@otter.itc.univie.ac.at  —  43 61 41 ext. 78

4.5 Group Description of GUP-Linz

General Information about the Group

• Group Name: ACPC Group at the University Linz

• Group Leader: Prof. Dr. J. Volkert

• Address:

Institute of Computer Science

Dept. for Graphics and Parallel Processing University of Linz Altenbergerstr. 69 A-4040 Linz, Austria, Europe

• Phone Number: ++43-732-2468-887 (secretary)

• FAX Number: ++43-732-2468-10

• Domain Name: Internet: gup.uni-linz.ac.at

• Persons:

Name  Internet-account Bitnet-account Phone Number
Jens Volkert  volkert  –  888
Siegfried Grabner  grabner  –  884
Dieter Kranzlmüller  dk  –  885
Alfred Spalt  spalt  –  885
Wolfgang Stürzlinger  wrzl  –  884

4.5.1 nCUBE 2 Model 1

Footprint

System  nCUBE 2
Architecture  Distributed Memory MIMD
# of Processors  64

4 IO Processors

Type of Processors nCUBE Custom Chip
Memory Size  4 MB per node
Disk Capacity  4 GB
Operating System Vertex, Vortex
Name  ncube00.edvz.uni-linz.ac.at
Internet Address  140.78.3.11
Location  UNI Linz
System Administrator S. Grabner
Network Coordinator S. Grabner

Availability

Group RISC TU-Vienna Vienna GUP Salzburg World

good

Further Description This machine is used for research and educational purpose. Footprint of the Host Computer

System  SUN4/330
Type of Processors SPARC
Memory Size  56 MB
Disk Capacity  1.4 GB
Operating System BSD 4.3
Name  nCube.edvz.uni-linz.ac.at
Internet Address  140.78.3.5
Location  UNI Linz
System Administrator S. Grabner
Network Coordinator S. Grabner

Software NCC

• Name: ncc

• Location: /usr/ncube/current/bin

• Type: Crosscompiler for nCUBE 2 programs on SUN330

• Characterization: compiles C, FORTRAN, and Assembler code for the nCUBE on the SUN frontend. ncc is used like the normal cc compiler.

• Functions:

– Invocation: ncc

– Termination:

       – Help: nman ncc
       – Example: ncc program.c

– Other important functions:

• Installation: include the path /usr/ncube/current/bin

• Manuals: nCUBE 2 Programmers Guide, Chapter 6 XNC

• Name: xnc

• Location: /usr/ncube/current/bin/sun4

• Type: Loader for nCUBE 2 programs down to nCUBE nodes.

• Characterization:

• Functions:

– Invocation: xnc

– Termination: nkill processID – to terminate node programs

    – Help: nman xnc

– Example: xnc -d 4 executable – start a program on 16 nodes

– Other important functions:

• Installation: include the path /usr/ncube/current/bin/sun4 or

  /usr/ncube/current/bin/`arch`, respectively.

• Manuals: nCUBE 2 Programmers Reference (= nman page for xnc)

NDB • Name: ndb

• Location: /usr/ncube/current/bin/sun4

• Type: Symbolic source code and assmebly level debugger for nCUBE 2.

• Characterization:

• Functions:

– Invocation: ndb

– Termination: quit

    – Help: nman ndb

– Example: ndb -p processID executable – if program is hung

– Note: the source code must be compiled with ncc and the -g option.

• Installation: include the path /usr/ncube/current/bin/sun4 or

  /usr/ncube/current/bin/`arch`, respectively.

• Manuals: nCUBE 2 Programmers Guide, Chapter 7 XTOOL

• Name: xtool

• Location: /usr/ncube/current/bin/sun4

• Type: Execution Profiler for nCUBE 2 programs.

• Characterization: analyzes the execution profile of FORTRAN and C programs (graphically).

• Functions:

– Invocation: xtool programname

– Termination:

    – Help: nman xtool or xtool "-?"

– Example:

– Other important functions:

• Installation: include the path /usr/ncube/current/bin/sun4 or

  /usr/ncube/current/bin/`arch`, respectively.

• Manuals: nCUBE 2 Programmers Reference, chapter 8

CTOOL • Name: ctool

• Location: /usr/ncube/current/bin/sun4

• Type: Communication Profiler for nCUBE 2 programs.

• Characterization: analyzes the communication profile of FORTRAN and C programs (graphically).

• Functions:

– Invocation: ctool options

– Termination:

    – Help: nman ctool or ctool "-?"

– Example:

– Other important functions:

• Installation: include the path /usr/ncube/current/bin/sun4 or

  /usr/ncube/current/bin/`arch`, respectively.

• Manuals: nCUBE 2 Programmers Reference, chapter 9

4.6 Group Description of the Institute for Software Technology and Parallel Systems, University of Vienna (PPS)

General Information about the Group

• Group Name: Research Group Prof. H.P. Zima

• Group Leader: Prof. H.P. Zima

• Address:

INSTITUT FÜR SOFTWARETECHNIK UND PARALLELE SYSTEME UNIVERSITÄT WIEN o.Univ.Prof.Dr. H. Zima Brünner Str. 72 A-1210 Vienna, AUSTRIA

• Phone Number: ++43-1-392647/221 (Head)

• Phone Number: ++43-1-392647/222 (Secretary)

• Phone Number: ++43-1-40103/2790 (ACPC coordinator)

• FAX Number: ++43-1-392647/224

• Internet domain name: par.univie.ac.at

• Bitnet domain name: awiuni11.bitnet

• Persons at Brünner Str. 72:

Name  Internet-account Bitnet-account Phone Number
Prof. Dr. H.P. Zima  zima  A4424DAJ  -221
M. Morawitz (sec.)  sec  -222
E. Obermaier (sec.)  sec  -222
DI. S. Benkner  sigi  -202
Dr. Ing. R. Blasko  roman  -226
Dr. P. Brezany  brezany  -227
B. Chapman B.Sc.(Hons.)  barbara  -205
DI. T. Fahringer  tf  -201
DI. H. Moritsch  hm  -202
Dr. A. Schwald  schwald  -201
Dr. V. Sipkova  sipka  -207
E. Cikan (tech.)  cikan  -206

• Persons at Rathausstr. 19/II/3:

Name  Internet-account Bitnet-account  Phone Number
DI. B. Knaus  bk  A4424DAI  ++43-1-40103-2790
DI. P. Kutschera  peter  A4424DAK ++43-1-40103-2789

4.6.1 Intel iPSC/860 Hypercube

Footprint

System  Intel iPSC/860
Architecture  MIMD Distributed Memory
# of Processors  16
Type of Processors i860
Memory Size  8MB per processor
Disk Capacity  1.3GB on CFS
Operating System NX/2
Name  cube
Internet Address  131.130.70.34
Location  Institut für Statistik und Informatik

Lehrkanzel Zima Brünner Str. 72 1210 Wien

System Administrator Barbara Chapman, barbara@par.univie.ac.at

Network Coordinator Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
y y y y y n

Further Description Footprint of the Host Computer

System  Frontend for Intel Hypercube
Type of Processors Intel 80386

Memory Size

Disk Capacity  380MB
Operating System UNIX System 5, V 3.2
Name  pixy
Internet Address  131.130.70.33
Location  Institut für Statistik und Informatik

Lehrkanzel Zima Brünner Str. 72 1210 Wien

System Administrator Barbara Chapman, barbara@par.univie.ac.at

Network Coordinator

Software

PICL

• Name: Portable Instrumented Communication Library, Version 1.92

• Location: /newpicl (documentation in /picl-org)

• Type: Library package for portable parallel programming

• Characterization: PICL provides a library of routines which may be used instead of the native communication routines on any machine on which it is installed. Timestamped trace data is generated, automatically or as specified by the user (communication, idle times, events).

• Software Description: PICL provides portable syntax for parallel programs across several multiprocessor systems, including the Intel hypercubes and simulator, the Delta, the NCUBE and Cogent systems. It contains routines for communication primitives, such as send and receive, and related system calls, and widely-used higher routines such as global summing and broadcasts.

It also has a built-in tracing facility, which may be controlled by the user. This enables the gathering of data on interprocessor communication, processor idle times, and user-defined events. A separate software package, ParaGraph, displays the tracing output graphically. This is currently installed on our local SUN network only. However, versions of ParaGraph are available for several types of workstation. Contact the system administrator for information on obtaining the source code.

The most convenient way to use the PICL routines is to include the appropriate library at compile time. Separate libraries for the host and node code have been generated and are to be found in the directory /newpicl. Examples of usage of the routines are to be found in the User’s Guide (see below).

Our installation comprises the library packages and the source code for the libraries in the subdirectories picl and port. Note that this version has been optimized for use on the iPSC/860. The original version of the PICL libraries is available in source form only in the directory /picl-org and its subdirectories.

• Installation:

The appropriate libraries should be linked in when the host and node programs are compiled. A description of how to do this is to be found in README-USAGE.

• Manuals: There are two manuals. The user’s guide is located in userguide.tex (and .sty). The C reference manual is provided in cref.tex. Both are located in the directory /picl-org. These document the original version of PICL. The modifications are described in the two upgrade files in /newpicl. Note that there is currently no Fortran reference manual.

4.6.2 Transputersystem

Footprint

System  Impuls 2400
Architecture  Transputersystem
# of Processors  16
Type of Processors T800
Memory Size  1 MByte
Disk Capacity  0
Operating System iserver
Name  —
Internet Address  —
Location  Rathausstr. 19/2/3

A-1010 Wien

System Administrator Peter Kutschera (peter@merlin.par.univie.ac.at)

Network Coordinator

Availability

Group RISC TU-Vienna PPS Vienna Linz Salzburg World
y  y  y  y  y  y  y

To get an account ask peter@merlin.par.univie.ac.at .

Further Description

System  HP9000/425s
Type of Processor 68040
Memory Size  16 MB
Disk Capacity  1.2 GB
Operating System Domain/OS 10.3.5
Name  merlin.par.univie.ac.at
Internet Address  131.130.34.65
Location  Rathausstr. 19/2/3

A-1010 Wien

System Administrator Peter Kutschera (peter@merlin.par.univie.ac.at)

Network Coordinator

The system is also connected to an PC (no access from the network).

Software C–Compiler

• Name: INMOS C compiler Version 2.01.03 (built at 17:51:52 Oct 16 1990)

• Location: icc and tools are in /usr/local/bin

• Type: Compiler, linker,…

• Characterization: Takes C-sources with library-calls for communication and a description of the configuration of the transputer–system and creates a bootable file. This file can be executed on the transputers.

• Functions: NO automatic parallelization or communication – Invocation: icc

– Termination: none

    – Help: manuals
    – Example: see /usr/local/ictools/examples

– Other important functions:

• Installation: All used environment-vars are automatically set to defaults.

• Manuals: ANSI C Toolset User Manual, INMOS Limited ANSI C Toolset Language Reference, INMOS Limited Occam2–Compiler

• Name: occam 2 compiler Version 2.01.01 (21:15:57 Feb 27 1991) (Transputer)

• Location: oc and tools are in /usr/local/bin

• Type: Compiler, linker,…

• Functions:

– Invocation: oc

– Termination: none

    – Help: Manuals
    – Example: see /usr/local/d7205/examples

– Other important functions:

• Installation: All used environment-vars are automatically set to defaults.

• Manuals: OCCAM User Manual Part 1: User Guide and Tools, INMOS Limited OCCAM User Manual Part 2: Occam Libraries, INMOS Limited

config • Name: config, Version 0.2

• Location: /usr/local/bin

• Type: Utilitiy

• Characterization: Tool to set or determine the configuration of the links between the transputers.

• Functions:

    – Invocation: config options
    – Help: config

– Example:

      config -net 4 4           # set link-switches
      config -show              # check configuration
      config -cr_cfs > test.cfs # create Hardware-Part of .cfs-file

• Installation: none

• Manuals: none

4.6.3 MasPar

Footprint

System  MasPar MP-1101A
Architecture  Massively parallel SIMD
# of Processors  Data Parallel Unit (DPU)

with 1024 Processor Elements (PE)

Type of Processors PE: full-custom processor chip (1.8 MIPS)
Memory Size  DPU: 16 KBytes in each PE, 1.128 MBytes in ACU
Disk Capacity  —
Operating System —
Name  —
Internet Address  —
Location  Institut fuer Statistik und Informatik

Lehrkanzel Zima

Bruenner Str. 72, 1210 Wien

System Administrator Viera Sipkova sipka@par.univie.ac.at

Network Coordinator

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World

Further Description

MasPar MP-1 is a massively parallel, data-parallel, fine-grained processing system. It consists of 3 Subsystems: Front-End (FE), Data Parallel Unit (DPU), and I/O Subsystem. The FE is a DEC 5000 workstation which runs UNIX operating system with graphics and network services, including the software environment in which all the MasPar tools and utilities are executed. The DPU performs all parallel processing. It is composed of the Array Control Unit (ACU), Processor Element Array (PE Array) and Communication Mechanisms. The ACU is a register- based (14 MIPS) processor with its own registers, data memory (128 KBytes), and instruction memory (1 MByte - 4 GBytes). The PE array consits of 1K PEs that are arranged in a 2-dimnesional matrix (32x32). Each PE contains the arithmetic- logical unit (1.8 MIPS processor), registers, and 16 KBytes of RAM. Communications are implemented via VME Bus, ACU-PE Bus, and X-Net and Global Router.

Peak performance of MasPar MP-1: Each PE operates at 1.6 MIPS (32-bit integer add) and 80Kflops (average of 32-bit floating point multiply and add). MP-1 systems currently support PE arrays with 1,024 to 16,384 processor elements. Peak performance ranges from 1,600 to 26,000 integer MIPS (32-bit add), from 82 to 1,300 single precision Mflops (average of add and multiply) and from 36 to 580 double precission (64-bit) Mflops.

Footprint of the Host Computer

System  DEC Station 5000

Type of Processors

Memory Size  24 MB
Disk Capacity  665 MByte hard disk

optical compact disc drive (635 MByte)

Operating System ULTRIX
Name  maspar
Internet Address  131.130.70.36
Location  Institut fuer Statistik und Informatik

Lehrkanzel Zima

Bruenner Str. 72, 1210 Wien

System Administrator Viera Sipkova sipka@par.univie.ac.at

Network Coordinator

Software

ULTRIX and Misc. Tools

• Name: ULTRIX Worksystem Software, Version 4.0 (ULTRIX Operating System; DEC-Windows System); MasPar System Software Version 2.1, Version 2.2

• Location:

• Type: Programming languages: MPL (MasPar Application Language), MPF (MasPar Fortran); Compilers: mpl, mpfortran, mpl cc, mpl fort; Programming environment MPPE.

• Characterization: The mpl language is based on K&R C and is used for programming the DPU. The MPF language is based on FORTRAN 77 with extensions from DEC-s VAX Fortran and the new Fortran 90 array statements. The MPPE is a set of tools for running, debugging and analysing of parallel programs.

• Functions: Compilers and the MPPE are used in a way which is standard for UNIX operating system and DEC-windows.

  – Invocation: in standard way for UNIX and DEC-windows
  – Termination: in standard way for UNIX and DEC-windows
  – Help: by online manuals

  – Example: mpfortran -o exam exam.f

– Other important functions:

• Installation: setenv MP P AT H/usr/maspar; setpathMP PATH/bin

• Manuals: MP-1 System Hardware manuals Version 2.1, MasPar Overview and MPPE manuals Version 2.1, Version 2.2 , MasPar MPL Manuals Version 2.1, Version 2.2, MasPar Fortran Programming Manuals Version 2.1, ULTRIX Manuals Version 4.0

Location: Institut fuer Statistik und Informatik, Bruenner Str. 72, 1210 Wien

4.7 Group Description of RIST++ Salzburg

General Information about the Group

• Group Name: Research Institute for Software Technology + PLUS (Paris Lodron University Salzburg) (RIST++ -Salzburg).

• Group Leader: Prof. Dr. Peter Zinterhof.

• Address:

RIST++

Research Institute for Software Technology Paris Lodron University Salzburg Hellbrunnerstr. 34 A-5020 Salzburg, Austria, Europe

RIST++

Research Institute for Software Technology Paris Lodron University Salzburg Jakob Haringerstr. 5 A-5020 Salzburg, Austria, Europe

• Phone Number: 0662-8044 (secretary: 0662-8044-6715).

• FAX Number: 0662-8044-5010

• Persons at Hellbrunnerstr. 34:

Domain Name: Internet: EDVZ.SBG.AC.AT

Name  Internet-account Bitnet-account Phone Number
Prof. Dr. P. Zinterhof  ZINTERHOF  6714
Christine Huber (sec.)  HUBERC  6715
Mag. Erika Hausenblas HAUSENBLAS  5308
Mag. Andrea Lauß  LAUSS  5323
Wolfgang Schmid  SCHMID  5308
Stefan Wegenkittl  5308
Peter Zinterhof jun.  ZINTER2  6727

• Persons at Jakob Haringerstr. 5:

Domain Name: Internet: kirk.coma.sbg.ac.at

Name  Internet-account Bitnet-account Phone Number
Ing. Ernst Forsthofer  forsthof  6770
Dr. Norbert Ulamec  ulamec  6770
Mag. Markus Salchegger  salchegg  6771
Peter Schmitzberger  schmitz  6771

4.7.1 SUN 3/60

Footprint

System  SUN 3/60
Architecture  Single Processor
# of Processors  1
Type of Processors 68020
Memory Size  12 MB
Disk Capacity  1 GB
Operating System SUN OS 4.1
Name  kirk.coma.sbg.ac.at
Internet Address  141.201.6.2
Location  RIST++,

Uni Salzburg

System Administrator E. Forsthofer
Network Coordinator E. Forsthofer

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
y  y  y  y  y  y

Software STRAND\underline{88}

• Name: STRAND\underline{88} Buckingham Release

• Location: /usr/local/strand

• Type: Parallel Declarative Programming Language

• Characterization: STRAND\underline{88} is an implementation of the Strand language, which includes an environment for developing and running Strand programs. The Strand language is desinged to utilize the facilities of concurrent machines and parallel architectures. It supports concurrent processing, stream communication between processes, Data-flow, modular development, declarative reading, object oriented programming and interfaces to other languages and data types.

• Functions:

– Invocation: st

– Termination: <CTRL>C

– Help: st -h – Example:

– Other important functions:

   • Installation: source /usr/local/strand/strandrc

• Manuals: STRAND\underline{88} User Manual, Book: Strand New Concepts in Parallel Programming

4.7.2 SPARC 2

Footprint

System  Sparc 2
Architecture  Single Processor
# of Processors  1
Type of Processors SPARC
Memory Size  32 MB
Disk Capacity  420 MB
Operating System SUN OS 4.1.1
Name  spock.coma.sbg.ac.at
Internet Address  141.201.6.3
Location  RIST++,

Uni Salzburg

System Administrator E. Forsthofer
Network Coordinator E. Forsthofer

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
y  y  y  y  y  y

Software STRAND\underline{88}

4.7.3 SPARC ELC

Footprint

System  Sparc ELC
Architecture  Single Processor
# of Processors  1
Type of Processors SPARC
Memory Size  8 MB
Disk Capacity  0 MB
Operating System SUN OS 4.1.1
Name  scotty.coma.sbg.ac.at
Internet Address  141.201.6.4
Location  RIST++,

Uni Salzburg

System Administrator E. Forsthofer
Network Coordinator E. Forsthofer

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
y  y  y  y  y  y

Software STRAND\underline{88}

4.7.4 CONVEX C220

Footprint

System  CONVEX C220
Architecture  Parallel Vector
# of Processors  2
Type of Processors Vector-processor
Memory Size  128 MB
Disk Capacity  8 GB
Operating System CONVEX OS 9.1
Name  CONVEX
Internet Address  141.201.10.2
Location  TECH Consult

Salzburg

System Administrator Kurt Krenn
Network Coordinator Kurt Krenn

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World

good

Further Description Software Vectorizing C-Compiler

• Name: CONVEX Vectorizing C-CompilerV 2.0

• Location:

• Type: Parallelizing and Vectorizing Compiler

• Characterization:

• Hints: To see, how to use the HIGH PERFORMANCE MATHEMATICAL VECTOR LIBRARY (VECLIB) type: man veclib. There are many vectorized functions (subroutines) you can use in your application-programs.

• Functions:

– Invocation: cc

– Termination: <CTRL>C

– Help: man cc

– Example:

– Other important functions:

• Installation:

• Manuals: Vectorizing C-Compiler User Manual,

4.7.5 ALPHA Cluster (6+1 DEC 3000 Alpha AXP 400)

Footprint

System  1× DEC 3000 Alpha AXP 400
Architecture  Single Processor
# of Processors  1
Type of Processors DECchip 21064/133 MHz
Memory Size  64 MB
Disk Capacity  1 GB
Operating System DEC OSF/1 T1.2-2
Name  picard.coma.sbg.ac.at
Internet Address  141.201.65.2
Location  RIST++,

Uni Salzburg

System Administrator E. Forsthofer
Network Coordinator E. Forsthofer

System  6× DEC 3000 Alpha AXP 400
Architecture  Single Processor
# of Processors  1
Type of Processors  DECchip 21064/133 MHz
Memory Size  32 MB
Disk Capacity  1 GB
Operating System  DEC OSF/1 T1.2-2

Names and addresses

Name  riker.coma.sbg.ac.at
Internet Address  141.201.65.6
Name  laforge.coma.sbg.ac.at
Internet Address  141.201.65.7
Name  worf.coma.sbg.ac.at
Internet Address  141.201.65.8
Name  crusher.coma.sbg.ac.at
Internet Address  141.201.65.9
Name  data.coma.sbg.ac.at
Internet Address  141.201.65.10
Name  troi.coma.sbg.ac.at
Internet Address  141.201.65.11
Location  RIST++,

Uni Salzburg

System Administrator E. Forsthofer
Network Coordinator E. Forsthofer

Availability

Group RISC TU-Vienna Vienna Linz Salzburg World
y y y y y y

Software DEC C++ for DEC OSF/1 AXP DEC FUSE for DEC OSF/1 AXP DEC Fortran for DEC OSF/1 AXP

Appendix A

Table of All Available Parallel Machines in the ACPC Network

This appendix can be seen as a summary of Chapter 4 as it shows all parallel machines in the ACPC Network in just one table. The information given in this table is sufficient to set up a connection with any machine in the ACPC Network (even if there is no name service available). However, if one wants to know details of the machines mentioned in Table A.1 one has to look for the corresponding section in Chapter 4.

As several machines like the Transputers and the nCube can only be used from a host machine, for these machines this host is put into Table A.1.

Group  Type of Machine Symbolic Address  Numeric Address
RISC-Linz Sequent Symmetry melmac.risc.uni-linz.ac.at 193.170.36.100
Transputer  venus.risc.uni-linz.ac.at^{1}  193.170.36.70
Uni-Vienna Transputer  epikur.ani.univie.ac.at^{1}  131.130.32.113
TU-Vienna Sequent Balance  eacpc1.tuwien.ac.at  128.130.36.3
Intel Hypercube  eacpc2.tuwien.ac.at^{1}  128.130.36.4
GUP-Linz nCube  nCube.edvz.uni-linz.ac.at^{1} 140.78.3.5
Uni-Vienna Intel Hypercube  pixy.par.univie.ac.at^{1}  131.130.70.33
(PPS)  MasPar  maspar.par.univie.ac.at^{1} 131.130.70.36
Transputer  merlin.par.univie.ac.at^{1}  131.130.34.65
RIST++  Alpha Cluster  picard.coma.sbg.ac.at  141.201.65.2
(Salzburg) Convex  convex.edvz.sbg.ac.at  141.201.10.2

Table A.1: List of All Parallel Machines in the ACPC Network

1This is the address of the host computer that has to be used to get access to the parallel machine.

Appendix B

Alphabetical List of Persons in the ACPC

This appendix summarizes the most important information about persons that are working in the ACPC by associating with each person her/his email address and telephone number. More information for each person can be found in the appropriate group description of Chapter 4.

*_{Index}

Benkner, S.

sigi@par.univie.ac.at 0222 392647 202

Blasko, R.

roman@par.univie.ac.at 0222 392647 226

Brezany, P.

brezany@par.univie.ac.at 0222 392647 227

Buchberger, Prof. Bruno buchberg@risc.uni-linz.ac.at 07236 3231 41

Chapman, B.

barbara@par.univie.ac.at 0222 392647 205

Cikan, E.

cikan@par.univie.ac.at 0222 392647 206

Fahringer, T.

tf@par.univie.ac.at 0222 392647 201

Ferscha, Alois ferscha@ani.univie.ac.at 0222 4086366 18

Forsthofer, Ernst forsthof@kirk.coma.sbg.ac.at 0662 8044 6770

Fritscher, Josef fritscher@edvz.tuwien.ac.at 0222 58801 5505

Grabner, Siegfried grabner@gup.uni-linz.ac.at 0732 2468 884 Haring, Prof. Günter

haring@ani.univie.ac.at

0222 4086366 12 Hausenblas, Erika

hausenblas@edvz.sbg.ac.at 0662 8044 5308

Hofacker, I.

ivo@otter.itc.univie.ac.at 0222 436141 78

Hong, Hoon hhong@risc.uni-linz.ac.at 07236 3231 48

Huber, Christine huber@edvz.sbg.ac.at 0662 8044 6715 Husinsky, Irmgard

husinsky@edvz.tuwien.ac.at 0222 58801 5484

Jebelean, Tudor tjebelea@risc.uni-linz.ac.at 07236 3231 64

Kleinert, Dr. Wolfgang kleinert@edvz.tuwien.ac.at 0222 58801 5480

Knaus, B.

bk@par.univie.ac.at 0222 40103 2790

Kotsis, Gabriele kotsis@ani.univie.ac.at 0222 4086366 14 Kranzlmüller, Dieter

dk@gup.uni-linz.ac.at

0732 2468 885 Krommer, Arnold Kutschera, P. peter@par.univie.ac.at 0222 40103 2789

Landerl, Werner landerl@risc.uni-linz.ac.at 07236 3231 67

Lauß, Andrea lauss@edvz.sbg.ac.at 0662 8044 5323

Loidl, Hans Wolfgang hwloidl@risc.uni-linz.ac.at 07236 3231 66

Morawitz, M.

sec@par.univie.ac.at 0222 392647 222

Moritsch, H.

hm@par.univie.ac.at 0222 392647 202

Obermaier, E.

sec@par.univie.ac.at 0222 392647 222

Pigel, Georg pigel@ani.univie.ac.at 0222 4086366 13

Presslauer, Ursula presslauer@ani.univie.ac.at 0222 4086366 23

Salchegger, Markus salchegg@spock.coma.sbg.ac.at 0662 8044 6771

Schmid, Wolfgang schmid@edvz.sbg.ac.at 0662 8044 5308

Schmitzberger, Peter schmitz@spock.coma.sbg.ac.at 0662 8044 6771

Schreiner, Wolfgang schreine@risc.uni-linz.ac.at 07236 3231 66

Schuster, Prof. P.

pks@tiger.itc.univie.ac.at 0222 436141 78 Schwald, A.

schwald@par.univie.ac.at 0222 392647 201

Siegl, Kurt siegl@risc.uni-linz.ac.at 07236 3231 67

Sipkova, V.

sipka@par.univie.ac.at

0222 392647 207 Spalt, Alfred

spalt@gup.uni-linz.ac.at 0732 2468 885 Stürzlinger, Wolfgang

wrzl@gup.uni-linz.ac.at

0732 2468 884

Teodosiu, Dan danteo@risc.uni-linz.ac.at

Überhuber, Dr. Christoph 0222 58801 5412

Ulamec, Norbert ulamec@kirk.coma.sbg.ac.at 0662 8044 6770

Volkert, Prof. Jens volkert@gup.uni-linz.ac.at 0732 2468 888

Wabnig, Harald wabnig@ani.univie.ac.at 0222 4086366 17

Wegenkittl, Stefan 0662 8044 5308

Zettler, Peter zettler@risc.uni-linz.ac.at 07236 3231 67 Zima, Prof. H. P.

zima@par.univie.ac.at

0222 392647 221 Zinterhof, Prof.

zinterhof@edvz.sbg.ac.at 0662 8044 6714 Zinterhof jun., Peter

zinter2@edvz.sbg.ac.at

0662 8044 6727